### v2

# Reinforcement Learning from Bagged Reward

Abstract
In Reinforcement Learning (RL), it is commonly assumed that an immediate reward signal is generated for each action taken by the agent, helping the agent maximize cumulative rewards to obtain the optimal policy. However, in many real-world scenarios, immediate reward signals are not obtainable; instead, agents receive a single reward that is contingent upon a partial sequence or a complete trajectory. In this work, we define this challenging problem as Reinforcement Learning from Bagged Reward (RLBR), where sequences of data are treated as bags with non-Markovian bagged rewards. We provide a theoretical study to establish the connection between RLBR and standard RL in Markov Decision Processes (MDPs). To effectively explore the reward distributions within these bags and enhance policy training, we propose a Transformer-based reward model, the Reward Bag Transformer, which employs a bidirectional attention mechanism to interpret contextual nuances and temporal dependencies within each bag. Our empirical evaluations reveal that the challenge intensifies as the bag length increases, leading to the performance degradation due to reduced informational granularity. Nevertheless, our approach consistently outperforms existing methods, demonstrating the least decline in efficacy across varying bag lengths and excelling in approximating the original MDP’s reward distribution. The code is available at an anonymous link: https://anonymous.4open.science/r/RLBR-F66E/.
11footnotetext: Equal contribution.
1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various domains, including autonomous driving Kiran et al. (2021), healthcare Yu et al. (2021), complex game playing Silver et al. (2016); Wurman et al. (2022), and financial trading Yang et al. (2020). One common and essential assumption for most RL algorithms is the availability of immediate reward feedback at each time step of the decision-making process. However, this assumption is violated in many real-world applications. Recognizing this gap, numerous studies Watkins (1989); Ke et al. (2018); Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023b) have explored the concept of delayed rewards, primarily focusing on trajectory feedback where rewards are allocated at the end of a sequence. Similarly, sparse reward settings, where agents receive infrequent and instance-specific feedback, present significant challenges in the well-known exploration-exploitation trade-off Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022).
On the other hand, real-world applications, such as autonomous driving (see Fig. 1), often feature complex non-immediate reward structures that neither sparse rewards nor trajectory feedback can fully capture. Providing reward for every action is impractical, and focusing only on end goals ignores crucial aspects of the journey. Typically, rewards are linked to completing specific tasks or sequences of actions, rather than individual actions or the final objective Early et al. (2022); Gaon and Brafman (2020). However, previous studies (see Fig. 2) mainly focused on learning desirable policies with immediate rewards Sutton and Barto (1998); Vinyals et al. (2019), trajectory feedback Watkins (1989); Arjona-Medina et al. (2019), or sparse rewards where feedback is infrequent and tied only to the current instance Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022), tending to fail under such scenarios.
To address these challenges, we introduce Reinforcement Learning from Bagged Rewards (RLBR), which better aligns with real-world scenarios by considering the cumulative effect of a series of state-action pairs. In RLBR, we define sequences of state-action pairs as bags, each associated with a cumulative bagged reward. This framework includes both the traditional RL setting, where each bag only contains a single instance (the first line of Fig. 2), and the trajectory feedback setting, where a bag spans an entire trajectory (the third line of Fig. 2). Furthermore, RLBR offers the potential to reduce the labeling workload by lessening the frequency of reward annotations. However, this benefit is balanced by increased learning complexity due to the reduced granularity of information.
Figure 1: An example of the reward bag structure on an autonomous driving trajectory. Each segment of the driving sequence is evaluated and assigned a score by an evaluator, and the score for each sequence is integrated based on the performance at each step.
In the RLBR framework, our focus is on leveraging bagged reward information to discern the significance of each instance within a bag and to understand the relationships among different bags. The challenge lies in accurately interpreting the contextual nuances within individual bags, as the instances within a bag are time-dependent on each other and their contributions to the bagged reward vary. Given the importance of context in RLBR, we turn to the bidirectional attention mechanism (Seo et al., 2016; Vaswani et al., 2017; Devlin et al., 2019), renowned for its effectiveness in contextual understanding, especially for time-dependent data. Specifically, we propose a Transformer-based reward model, leveraging the bidirectional attention mechanism to adeptly interpret context within a bag and allocate rewards to each instance accurately. This model can be utilized to enhance general RL algorithms, such as Soft Actor-Critic (SAC) Haarnoja et al. (2018), for environments with bagged rewards.
Our research contributes to the field in several ways. First, we establish the RLBR framework as a general problem setting and connect it theoretically to traditional Markov Decision Processes (MDPs), elaborated in Section 3. In Section 4, we introduce a Transformer-based reward model designed to assign rewards to individual instances while capturing environmental dynamics. Additionally, we propose an algorithm that alternates between optimizing this reward model and the policy, thereby enhancing the effectiveness of both components. Finally, in Section 5, our experiments show that the performance of baseline methods drops as the bag length increases, indicating that the larger bag length will increase the learning difficulty. Furthermore, we experimentally demonstrate the superiority of our method through comparative performance analyses and validate the ability of the proposed model to mimic the reward distribution of the ground truth MDP, highlighting its contextual understanding and adaptability to environmental dynamics.
2 Related Work
RL with Trajectory Feedback. RL with Trajectory Feedback (RLTF), termed episodic rewards or delayed rewards in some works (see the third line in Fig. 2), has become increasingly prominent in many applications (Watkins, 1989; Arjona-Medina et al., 2019; Zhang et al., 2023b). A key approach to this challenge is reward redistribution, aiming to assign rewards to individual instances more effectively. Return Decomposition for Delayed Rewards (RUDDER) Arjona-Medina et al. (2019) used a return-equivalent formulation for precise credit assignment, with advancements incorporating expert demonstrations and language models Patil et al. (2022); Liu et al. (2019); Widrich et al. (2021). Iterative Relative Credit Refinement (IRCR) Gangwani et al. (2020) presented a uniform reward redistribution model based on equal contributions from all state-action pairs, while Randomized Return Decomposition (RRD) Ren et al. (2021) proposed a novel upper bound for return-equivalent assumptions, integrating return decomposition with uniform reward redistribution. Additionally, Han et al. (2022) modified RL algorithms to use sequence-level information, helping agents learn from broader structures and long-term outcomes. However, our focus is on the original, unobservable rewards within reward bags, not the aggregated bagged rewards. Our experiments (see Section 5) show that the method by Han et al. (2022) is ineffective for long sequences.
Following the previous methodology, we adopt a reward redistribution learning strategy to enhance policy learning in the context of reward bags. However, conventional reward redistribution methods fail to effectively extract information from bag-level rewards, as they did not take the structure of the bag into consideration. Our experimental results demonstrate that directly applying previous methods to the reward bag setting does not yield results as promising as those achieved by our algorithm.
Figure 2: Comparison of four different types of reward settings in RL. In traditional RL, a reward is given based on each instance Sutton and Barto (1998); Vinyals et al. (2019), whereas in sparse reward settings, only some instances receive rewards Riedmiller et al. (2018); Devidze et al. (2022); Rengarajan et al. (2022). Both types of rewards are Markovian. Trajectory feedback and the proposed bagged reward both address non-Markovian situations; however, trajectory feedback provides only one reward signal for the entire trajectory Watkins (1989); Arjona-Medina et al. (2019), while bagged rewards can include multiple reward signals within the sequence.
RL with Sparse Rewards. The sparse reward setting (see the second line in Fig. 2) presents significant challenges due to infrequent feedback, making it difficult for agents to effectively explore the environment and discover successful strategies. To address this challenge, various methods have been developed to enhance exploration. Reward shaping strategies Ng et al. (1999); Hu et al. (2020); Tambwekar et al. (2019) added rewards to actions in a way that guides the agent towards better policies without altering the original reward function. Curiosity-driven methods Pathak et al. (2017); Sekar et al. (2020) encouraged agents to explore the environment by visiting unseen states, potentially solving tasks with sparse rewards. Additionally, curriculum learning in RL Florensa et al. (2018); Riedmiller et al. (2018) involved presenting an agent with a sequence of tasks with gradually increasing complexity, allowing the agent to eventually solve the initially given sparse reward task.
In contrast to sparse rewards, which are typically infrequent Markovian signals associated with individual state-action pairs, bagged rewards depend on the cumulative effect of sequences of state-action pairs and are non-Markovian, requiring context from the entire sequence to accurately assess the contribution of each instance.
Transformers for RL. Transformers, introduced by Vaswani et al. (2017), have proven effective in RL environments requiring high sample efficiency and generalization, such as StarCraft Vinyals et al. (2019); Zambaldi et al. (2019) and DMLab30 Parisotto et al. (2020). They have also been used for sequential modeling in offline RL Chen et al. (2021); Janner et al. (2021b) and as reward models in policy learning from offline datasets Kim et al. (2023). Luo et al. (2021) combined deep convolution transfer learning models and inverse RL for reward function acquisition, while Zhang et al. (2023a) transformed non-Markovian reward processes into Markovian ones, enhancing online interaction efficiency. For reward bags, where rewards are linked to sequences of state-action pairs, Transformers are ideal due to their ability to handle long-range dependencies. Our approach leverages a Transformer-based model to redistribute bag-level rewards into instance-level rewards, capitalizing on the Transformer’s proficiency in sequential data analysis and attention mechanisms.
3 Reinforcement Learning from Bagged Reward
In this section, we first provide preliminaries on RL with immediate rewards and trajectory feedback, and then formulate the RLBR problem with an extension of the traditional MDP, the Bagged Reward MDP (BRMDP). We further conduct a theoretical analysis of the relationship between the BRMDP and the traditional MDP, showing that solving RLBR problems not only relaxes the demands on reward acquisition but also guarantees finding the optimal policy even with limited reward information.
3.1 Preliminaries
We consider finite-horizon RL in this paper, which is traditionally modeled using a finite MDP, where rewards are promptly provided for each state-action pair (Sutton and Barto, 1998). This paradigm is encapsulated in a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathit{P},\mathit{r},\mu)$, with $s\in\mathcal{S}$ and $a\in\mathcal{A}$ as the sets of states and actions, $\mathit{P}$ as the state transition probability function, $\mathit{r}$ as the immediate reward function, and $\mu$ as the initial state distribution. The primary objective in this framework is to discover a policy $\pi=p(a|s)$ that maximizes the cumulative sum of rewards over a horizon length $T$:
$J(\pi)=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\bigg{]}.$ (1)
Distinct from this traditional approach, RLTF offers feedback only after a complete sequence of actions, or a trajectory (Arjona-Medina et al., 2019; Zhang et al., 2023b). A trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T-1},a_{T-1})\}$ includes $T$ state-action pairs, with a reward $R_{\mathrm{traj}}(\tau)$ that is the sum of latent immediate rewards $\sum_{t=0}^{T-1}r(s_{t},a_{t})$, but only the cumulative reward is observable at the end of the trajectory. Denoting by $\mathcal{T}(\pi)$ as the distribution of trajectories induced by $\pi,P,\mu$, the learning objective in RLTF is to maximize the expected trajectory-based reward:
$J_{\mathrm{traj}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\big{[}R_{\mathrm{ep}}(\tau)\big{]}.$ (2)
3.2 Problem Formulation
We formally define the setting of RLBR, which has a granularity of rewards in between the above two settings. First, we define the notion of bags, which are sub-pieces of complete trajectories. A trajectory $\tau$ is divided into several neighboring bags, and a bag of size $n_{i}$, which starts from time $i$, is defined as $B_{i,n_{i}}=\{(s_{i},a_{i}),\ldots,(s_{i+n_{i}-1},a_{i+n_{i}-1})\},0\leq i\leq i+n_{i}-1\leq T-1$. Afterward, we define the BRMDP to navigate the complexities of the aggregated non-Markovian reward signals:
Definition 3.1 (BRMDP).
A BRMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathit{P},\mathit{R},\mu)$, where
  • $\mathcal{S}$ and $\mathcal{A}$ are sets of states and actions.
  • $\mathit{P}$ is the state transition probability function.
  • $\mathit{R}$ denotes the reward function over bagged reward: $\mathit{R}(B_{i,n_{i}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})$.
  • $\mu$ represents the initial state distribution.
The essential properties of BRMDP are two-fold: 1) only bagged rewards can be obtained by the learner, not immediate rewards; 2) the bagged reward is the accumulated sum of immediate rewards within the bag. In the RLBR framework, a bag $B_{i,n_{i}}$ metaphorically aggregates individual rewards from a contiguous sequence of state-action pairs into a unified reward unit. A trajectory $\tau$ is a composite of a set of bags, denoted as $\mathcal{B}_{\tau}$, which ensures that each trajectory includes at least one reward bag. We further assume a bag partition function defined by the environment: $\mathcal{G}:\tau\rightarrow\mathcal{B}_{\tau}$, which is a task-dependent function for generating bags given an input trajectory. Consequently, the learning objective of the policy is to maximize the accumulated bagged rewards:
$J_{\mathrm{B}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\bigg{[}\sum_{B\in\mathcal{B}_{\tau}}R(B)\bigg{|}\mathcal{G}\bigg{]}.$ (3)
Notably, if each bag comprises a single instance ($n_{i}=1,\forall 0\leq i\leq T-1$), RLBR simplifies to standard RL. Conversely, if a single bag encompasses the entire trajectory ($n_{0}=T$), RLBR reduces to RLTF. This adaptability highlights the capacity of the RLBR framework to accommodate varying reward structure scenarios.
3.3 Equivalence of Optimal Policies in the BRMDP and the Original MDP
In the BRMDP, agents are unable to directly observe the immediate reward linked to each state-action pair. Nonetheless, we aim for agents to optimize policies in environments where bagged rewards are accessible and to perform well on the original task. To unify the learning objectives of the BRMDP with those of the traditional MDP, we formulate a theorem that crystallizes this relationship. The significance of this theorem lies in its assertion that the optimal policy derived in a BRMDP setting is congruent with that of an MDP when the bagged rewards are considered as cumulative sums. This finding not only validates the theoretical underpinnings of our approach but also affirms its practical relevance across various RL paradigms.
Theorem 1.
Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag. In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
The detailed proof of Theorem 1 is deferred to Appendix B. By this theorem, we can see that optimizing a policy on a BRMDP is equivalent to optimizing on the original MDP. The crux of the matter is how to accurately redistribute the bagged rewards so that the policies learned under these distributed rewards are more close to the policies on the original MDP.
Figure 3: The illustration of the Reward Bag Transformer (RBT) architecture. The Causal Transformer is used for reward representation by processing sequences of input data consisting of state-action pairs. The bidirectional attention layer is used for reward redistribution, utilizing the outputs of the Causal Transformer to predict instance-level rewards. Predicting the next state helps the model understand the environment, thereby improving reward prediction.
4 Reward Bag Transformer
This section delves into the intricacies of leveraging the bidirectional attention mechanism Seo et al. (2016); Vaswani et al. (2017); Devlin et al. (2019) for reward redistribution, designs a Transformer architecture, and outlines a comprehensive algorithm for the cultivation of efficient policies under the BRMDP framework.
4.1 Reward Redistribution based on Bidirectional Attention Mechanism
Building on the foundation of Theorem 1 that optimal policies in the BRMDP and MDP are equivalent, our focus shifts to the crucial process of redistributing rewards within a bag from the BRMDP. To capture the contextual influence of each instance within the sequence, the Causal Transformer Vaswani et al. (2017) is a natural choice as a sequential prediction model. Traditionally, Transformers in RL are used in a unidirectional manner Chen et al. (2021); Janner et al. (2021a); Micheli et al. (2023), where only previous instances influence the current prediction due to the unobservability of future instances. However, given that our bagged rewards are non-Markovian and both preceding and subsequent instances affect the contribution of the current instance to the bagged reward, a bidirectional attention mechanism Seo et al. (2016); Vaswani et al. (2017); Devlin et al. (2019) becomes pivotal. This mechanism connects both past and future instances within a bag, enabling a more comprehensive understanding of contextual influences. By quantitatively evaluating the contribution of each instance, the bidirectional attention mechanism facilitates nuanced and effective reward redistribution.
4.2 Proposed Approach
We introduce the Reward Bag Transformer (RBT), a novel approach designed for the BRMDP framework. The RBT is engineered to comprehend the complex dynamics of the environment through bags and to precisely predict instance-level rewards, facilitating effective reward redistribution.
Causal Transformer for Reward Representation. Referring to Fig. 3, the RBT comprises a Causal Transformer (Vaswani et al., 2017; Radford et al., 2018), which maintains the chronological order of state-action pairs Chen et al. (2021); Janner et al. (2021a). For each time step $t$ in a sequence of $M$ time steps, the Causal Transformer, represented as a function $f$, processes the input sequence $\sigma=\{s_{0},a_{0},\ldots,s_{M-1},a_{M-1}\}$, generating the output $\{x_{t}\}_{t=0}^{M-1}=f(\sigma)$. By aligning the output head $x_{t}$ with the action token $a_{t}$, we directly model the consequence of actions, which are pivotal in computing immediate rewards and predicting subsequent states, which in turn aids the model in better understanding the environmental dynamics.
Bidirectional Attention Layer for Reward Redistribution. Once we have obtained the output embeddings $\{x_{t}\}_{t=0}^{M-1}$, for reward prediction, they pass through a bidirectional attention layer to produce $\{\hat{r}_{t}\}_{t=0}^{M-1}$, where $\hat{r}_{t}\equiv\hat{r}_{\theta}(s_{t},a_{t})$ with $\theta$ being the RBT parameters. This layer addresses the unidirectional limitation of the Causal Transformer architecture Vaswani et al. (2017); Radford et al. (2018), integrating past and future contexts for enhanced reward prediction accuracy. For state prediction, $x_{t}$ is input into a state linear decoder, yielding the predicted next state $\hat{s}_{t+1}\equiv\hat{s}_{\theta}(s_{t},a_{t})$.
The core of the RBT architecture is its bidirectional attention mechanism. For each output embedding $x_{t}$, we apply three different linear transformations to obtain embeddings for query $\mathbf{q}_{t}\in\mathbb{R}^{d}$, key $\mathbf{k}_{t}\in\mathbb{R}^{d}$, and value $v_{t}\in\mathbb{R}$. Then the instance-level reward is calculated by
$\hat{r}_{t}=\sum_{\ell=0}^{M-1}\texttt{softmax}(\frac{\{\langle\mathbf{q}_{t},\mathbf{k}_{t^{\prime}}\rangle\}_{t^{\prime}=0}^{M-1}}{\sqrt{d}})_{\ell}\cdot v_{\ell},$ (4)
where $d$ is the embedding dimension of the key. The rescaling operation is used to prevent extremely small gradients as in Vaswani et al. (2017). This mechanism enables the RBT to consider both the immediate and contextual relevance of each state-action pair in the trajectory when predicting rewards.
Algorithm 1 Policy Optimization with RBT
1:  Initialize replay buffer $\mathcal{D}$, RBT parameters $\theta$.
2:  for trajectory $\tau$ collected from the environment do
3:     Store trajectory $\tau$ with bag information $\{(B_{i,n_{i}},R(B_{i,n_{i}}))\}_{B_{i,n_{i}}\in\mathcal{B}_{\tau}}$ in $\mathcal{D}$.
4:     Sample batches from $\mathcal{D}$ and estimate bag loss based on Eq. (7).
5:     Update RBT parameters $\theta$ based on the estimated loss.
6:     Relabel rewards in $\mathcal{D}$ using the updated RBT.
7:     Optimize policy using the relabeled data by off-the-shelf RL algorithms (such as SAC).
8:  end for
4.3 Learning Objectives
The learning objectives of the RBT are twofold: reward prediction within each reward bag and state transition forecasting. These objectives are critical for enabling the model to navigate the complex dynamics of BRMDP environments.
Reward Prediction. The RBT is trained to ensure that, for each reward bag, the sum of predicted instance-level rewards matches the total bagged reward. This is vital for maintaining the integrity of the reward structure in the BRMDP framework. The loss function for this objective is expressed as
$\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})=\left(\sum_{t=i}^{i+n_{i}-1}\hat{r}_{t}-R(B_{i,n_{i}})\right)^{2},$ (5)
where $B_{i,n_{i}}$ represents the reward bag starting at time step $i$ with length $n_{i}$, and $R(B_{i,n_{i}})$ is the total reward for this bag. This formulation encourages the RBT to learn a nuanced distribution of rewards across states and actions within a bag. At the same time, it ensures that the sum of redistributed rewards matches the total bagged reward, maintaining consistency as per Theorem 1.
State Transition Forecasting. Alongside reward prediction, the RBT is tasked with accurately predicting the next state in the environment given the current state and action. This capability is crucial for understanding the dynamics of the environment. The corresponding loss function is:
$\mathcal{L}_{\mathrm{s}}(\theta;s_{t},a_{t},s_{t+1})=\|\hat{s}_{t+1}-s_{t+1}\|^{2},$ (6)
[…]
Composite Loss. The final learning objective combines the reward and state prediction losses:
$\mathcal{L}_{\mathrm{bag}}$ $(\theta)=\underset{\tau\sim\mathcal{D}}{\mathbb{E}}\big{[}\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})\big{|}B_{i,n_{i}}\in\mathcal{B}_{\tau}\big{]}$ (7)
$+\beta\underset{\tau\sim\mathcal{D}}{\mathbb{E}}\big{[}\mathcal{L}_{\mathrm{s}}(\theta;s_{t},a_{t},s_{t+1})\big{|}(s_{t},a_{t},s_{t+1})\in\tau\big{]},$
where the coefficient $\beta>0$ balances the two loss components, and $\mathcal{D}$ denotes the replay buffer.
The RBT’s dual predictive capacity is its key advantage, enabling precise reward redistribution to individual instances and forecasting the next state. This leverages environmental dynamics for enhanced reward distribution as experimentally evidenced in the ablation study in Appendix D.2. Integrated with off-the-shelf RL algorithms such as SAC Haarnoja et al. (2018), the RBT can enhance policy learning within the BRMDP framework, as outlined in Algorithm 1.
5 Experiment
In the following experimental section, we scrutinize the efficacy of our proposed method using benchmark tasks from both the MuJoCo Brockman et al. (2016) and the DeepMind Control Suite Tassa et al. (2018) environments, focusing on scenarios with bagged rewards. We assess the performance of our method to understand its overall effectiveness and examine whether the RBT reward model accurately predicts rewards.
5.1 Compare with SOTA Methods
Experiment Setting. We evaluated our method on eight benchmark tasks from both the MuJoCo locomotion suite, including Ant-v2, Hopper-v2, HalfCheetah-v2, and Walker2d-v2, and the DeepMind Control Suite, including cheetah-run, quadruped-walk, fish-upright, and cartpole-swingup. Differing from standard environments where rewards are assigned at each step, our approach involved assigning a cumulative reward at the end of each bag while assigning a reward of zero to all other state-action pairs within the bag. The maximum length for each episode was fixed at 1000 steps across all tasks.
Baselines. In the comparative analysis, our framework was rigorously evaluated against several leading algorithms in the domain of RL with delayed reward:
  • SAC (Haarnoja et al., 2018): It directly utilized the original bagged reward information for policy training using the SAC algorithm.
  • IRCR Gangwani et al. (2020): It adopted a non-parametric uniform reward redistribution approach by using the sum of immediate per-step rewards as a stand-in for trajectory returns. We have adapted IRCR for bagged reward setting.
  • RRD Ren et al. (2021): It employed a reward model trained with a randomized return decomposition loss. We have adapted RRD for bagged reward setting.
  • LIRPG Zheng et al. (2018): It learned an intrinsic reward function to complement sparse environmental feedback, training policies to maximize combined extrinsic and intrinsic rewards. We use the same code provided by the paper.
  • HC Han et al. (2022): The HC-decomposition framework was utilized to train the policy using a value function that operates on sequences of data. This approach decoupled the value function approximation task for the current step from the historical trajectory. We employed the code as provided by the original paper.
While methods like RUDDER Arjona-Medina et al. (2019) and Align-RUDDER Patil et al. (2022) are known for addressing the problem of trajectory feedback, previous studies Gangwani et al. (2020); Ren et al. (2021); Zhang et al. (2023b) have shown superior performance using referenced methods. Additionally, since Align-RUDDER relies on successful trajectories for scoring state-action pairs, which is impractical in MuJoCo Patil et al. (2022), we ultimately excluded both methods from our comparison. Besides, detailed descriptions of the model parameters and hyper-parameters used during training are provided in Appendix C. More experimental results are included in Appendix D.1.
Figure 4: Performance comparison in MuJoCo (top row) and DeepMind Control Suite (bottom row) environments with six different fixed-length reward bag settings (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$). The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
Table 1: Performance comparison across arbitrary reward bag configurations over 6 trials with 1e6 time steps for training, presenting average scores and standard deviations. “Narrow” refers to bags with lengths varying arbitrarily from 25 to 200 and narrow intervals between -10 to 10. “Broad” denotes the setting with bag lengths varying arbitrarily from 100 to 200 and broad interval variations from -40 to 40. The best and comparable methods based on the paired t-test at the significance level $5\%$ are highlighted in boldface.
Bag Setting Environment SAC IRCR RRD LIRPG HC RBT(ours)
Narrow Ant-v2 0.87 368.69 2272.39 -756.78 106.92 5122.50
(2.98) (119.74) (835.86) (763.66) (153.86) (206.44)
Hopper-v2 317.72 3353.35 2184.41 126.13 510.66 3499.54
(52.17) (61.97) (807.71) (30.18) (94.49) (76.62)
HalfCheetah-v2 788.45 10853.85 9709.62 1101.38 4027.25 11282.24
(1737.57) (573.72) (1479.73) (1248.45) (441.01) (266.08)
Walker2d-v2 193.07 4144.65 3536.90 123.43 309.19 4983.39
(48.40) (673.66) (546.66) (50.97) (171.69) (311.09)
Broad Ant-v2 -3.31 368.69 1323.50 -1264.08 5.97 5167.79
(4.15) (158.46) (1079.60) (416.86) (20.08) (303.83)
Hopper-v2 329.48 3296.20 1102.38 203.01 701.84 3499.53
(44.21) (216.35) (892.12) (177.80) (149.44) (94.00)
HalfCheetah-v2 43.96 9158.14 4199.16 924.26 4460.80 10837.15
(94.32) (1402.62) (1476.85) (1110.97) (518.94) (254.99)
Walker2d-v2 176.09 4179.08 330.96 194.95 447.45 5202.38
(49.81) (937.42) (79.26) (98.05) (155.63) (248.35)
5.1.1 Fixed-Length Reward Bags
In the fixed-length reward bag experiment, we conducted experiments with six bag lengths (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$) across each environment. The experimental design followed the problem setting, which assumed that the conclusion of one reward bag directly preceded the beginning of the next. It aimed to illustrate the influence of varying bag lengths on the results, providing insight into how bag size affected the performance of the learning algorithm within these environments.
As shown in Fig. 4, the SAC method, using bagged rewards directly from the environment, suffers from a lack of guidance in agent training due to missing reward information. This issue worsens with longer bag lengths, indicating that increased reward sparsity leads to less effective policy optimization. The IRCR and RRD methods, treating rewards uniformly within a reward bag, outperform SAC, suggesting benefits from even approximate guidance. However, notable variance in their results indicates potential consistency and reliability issues. The LIRPG exhibits subdued performance across tasks, as it is proposed to solve sparse reward problems (Zheng et al., 2018), which is Markovian and does not align the reward bag setting. The HC method excels only with shorter bag lengths, suggesting that this value function modification method struggles to utilize information from longer sequences. The proposed RBT method consistently outperforms the other approaches across all the environments and bag lengths, showing that it is not only well-suited for environments with short reward bags but also capable of handling large reward bag scenarios. This demonstrated the capability of RBT to learn from the sequence of instances and, by integrating bagged reward information, accurately allocate rewards to instances, thereby guiding better policy training.
5.1.2 Arbitrary Reward Bags
To validate the effectiveness of our approach under more complex conditions, we designed an experiment that allowed for overlaps or gaps between reward bags, and the length of each bag was no longer fixed. This setup simulated more realistic scenarios and tested the robustness of our method.
The results, as detailed in Table 1, affirm the superior performance of the proposed RBT method in these complex reward settings. Notably, other baseline methods that rely on reward model training experience a significant drop in effectiveness, primarily due to sample scarcity impacting model accuracy and policy learning. In contrast, the IRCR method, which distributes rewards uniformly without a model, maintains its efficacy in some environments. This outcome suggests that while approximate rewards can still guide policy learning, incorrect rewards can be highly detrimental. The consistent success of our approach in various reward bag scenarios indicates its potential for application to a broader range of reward structures, highlighting its versatility and robustness in handling more intricate reward dynamics.
5.2 Case Study
Figure 5: Comparison of predicted rewards with true rewards and aggregated bagged rewards.
The previous experimental results showcase the superiority of RBT over baselines. This led to an intriguing inquiry: Is the RBT reward model proficient in accurately redistributing rewards? To investigate this question, we performed an experiment focused on reward comparison, utilizing a trajectory generated by an agent trained in the Hopper-v2 environment with a bag length of 100. As shown in Fig. 5, which spans 1000 steps, RBT-predicted rewards, unobservable true rewards, and observable bagged rewards (presented in a uniform format for better visualization) are compared. The figure indicates that the rewards predicted by the RBT closely match the trends of the true rewards. This observation suggest that the RBT is effective at reconstructing true rewards from bagged rewards, despite the coarse nature of the environmental reward signals.
6 Conclusion
In this paper, we introduce a general learning framework, Learning from Bagged Rewards (RLBR), and make theoretical connections between our learning objectives and traditional MDPs to ensure the justification of our approach. Building on this problem, we propose a Transformer-based reward model, the Reward Bag Transformer (RBT), to efficiently redistribute rewards by interpreting contextual information within bags as well as understanding environmental dynamics. The efficacy of RBT is demonstrated through extensive experiments, outperforming existing delayed-reward methods in various reward bag scenarios. Besides, our case studies highlight the RBT’s ability to effectively reallocate rewards, maintaining fidelity to the original MDP structure.
While the sum-form bagged reward currently integrated into RLBR addresses most scenarios, as evidenced by previous studies on trajectory feedback (Ren et al., 2021; Arjona-Medina et al., 2019; Zheng et al., 2018), there is value in exploring other forms of reward aggregation. Future research will investigate alternative reward structures, such as maximum values or complex combinations of latent rewards, to better capture the nuances of dynamic real-world environments. This exploration aims to enhance the adaptability and effectiveness of the RLBR framework in a broader range of applications.
Appendix A Broader Impact
In this work, we introduce the problem of Reinforcement Learning from Bagged Reward (RLBR), and propose the Reward Bag Transformer to address it. On the one hand, we recognize that these techniques could raise some potential issues. As collecting bagged rewards is much more convenient and natural than gathering instance-level rewards, this could lead to some risks of abusing unauthorized data. On the other hand, we believe that developing these techniques is still necessary for the importance of solving reinforcement learning tasks with bag-level feedback. Furthermore, there are many techniques for preserving data privacy, which can be compatible with our approach to avoid such problems.
Appendix B Omitted Proof
Theorem 1.
Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag. In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
Proof.
Let us consider the structure of bagged rewards within BRMDP. For a given bag $\mathit{{B}_{i,n_{i}}}$, the bagged reward $R(\mathit{{B}_{i,n_{i}}})$ is the sum of the individual rewards of the state-action pairs contained within it, given by:
$R(\mathit{{B}_{i,n_{i}}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t}).$
Over a complete trajectory $\tau$, the cumulative reward in BRMDP can be expressed as the sum of the rewards from all the bags along the trajectory:
$\sum_{B\in\mathcal{B}_{\tau}}R(B)=\sum_{i\in\mathcal{I}_{\tau}}R(\mathit{{B}_{i,n_{i}}})=\sum_{i\in\mathcal{I}_{\tau}}\bigg{(}\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})\bigg{)}=\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
where $\mathcal{I}_{\tau}$ denotes the set of initial timestep indices for $B\in\mathcal{B}_{\tau}$.
The policy optimization objective in BRMDP, $J_{\mathrm{B}}(\pi)$, aims to maximize the expected sum of bagged rewards along the trajectory from $t=0$, which is equivalent to maximizing the standard cumulative reward in MDP, $J(\pi)$:
$J_{\mathrm{B}}(\pi)$ $=\mathbb{E}_{\mathcal{T}(\pi)}\bigg{[}\sum_{B\in\mathcal{B}_{\tau}}R(B)\bigg{|}\mathcal{G}\bigg{]}$
$=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\bigg{]}=J(\pi).$
Given that the expected cumulative rewards for any policy in the BRMDP and MDP frameworks are equivalent, under the condition of infinite exploration or exhaustive sampling within the state-action space, the sets of optimal policies for each framework also coincide, implying that $\Pi=\Pi_{\mathrm{B}}$. ∎
Appendix C Experiment Settings and Implementation Details
Benchmarks with Bagged Rewards.
We introduced a novel problem setting in the suite of MuJoCo and DeepMind Control Suite locomotion benchmark tasks, termed as bagged rewards. Our simulations ran on the OpenAI Gym platform Brockman et al. (2016) and the DeepMind Control Suite Tassa et al. (2018), featuring tasks that stretched over long horizons with a set maximum trajectory length of $T=1000$. We used MuJoCo version 2.0 for our simulations, which is available at http://www.mujoco.org/. MuJoCo is licensed under a commercial license, and we have adhered to its terms of service and licensing agreements as stated on the official website. The DeepMind Control Suite is available under an Apache License 2.0, and we have complied with its terms of use.
In this setting, the agent received a bagged reward at the end of each collected bag. Reward bag experiments of different bag sizes (5, 25, 50, 100, 200, and 500) and trajectory feedback were set up to verify the effectiveness of the method. To evaluate the efficacy of proposed method, commonly used trajectory feedback algorithms were adapted to fit the bagged reward setting as baselines. In these experiments, each reward bag was treated as an individual trajectory, and these modified algorithms were applied accordingly. Additionally, experiments using standard trajectory feedback were conducted to provide a comparative baseline within the unique setting. The total episodic feedback was computed at the end of the trajectory and was the sum of the per-step rewards the agent had collected throughout the episode. This experiment setting was the same as some previous works for learning from trajectory feedback Gangwani et al. (2020); Ren et al. (2021).
Implementation Details and Hyper-parameter Configuration.
Table 2: Hyper-parameters of RBT.
Hyper-parameter Value
Number of Causal Transformer layers 3
Number of bidirectional attention layers 1
Number of attention heads 4
Embedding dimension 256
Batch size 64
Dropout rate 0.1
Learning rate 0.0001
Optimizer AdamW Loshchilov and Hutter (2018)
Weight decay 0.0001
Warmup steps 100
Total gradient steps 10000
In our experiments, the policy optimization module was implemented based on soft actor-critic (SAC) Haarnoja et al. (2018). We evaluated the performance of our proposed methods with the same configuration of hyper-parameters in all environments. The back-end SAC followed the JaxRL implementation Kostrikov (2021), which is available under the MIT License.
The RBT reward model was developed based on the GPT implementation in JAX Frostig et al. (2018), which is available under the Apache License 2.0. Our experiments utilized the Causal Transformer with three layers and four self-attention heads, followed by a bidirectional self-attention layer with one self-attention head. For detailed hyper-parameter settings of the RBT, please refer to Table 2.
For the baseline methods, the IRCR Gangwani et al. (2020) method was implemented based on the descriptions provided in the original paper. The RRD Ren et al. (2021) and LIRPG Zheng et al. (2018) methods are both licensed under the MIT License. The code of HC Han et al. (2022) is available in the supplementary material at https://openreview.net/forum?id=nsjkNB2oKsQ.
To ensure uniformity in the policy optimization process across all methodologies, each was subjected to 1,000,000 training iterations. For the proposed method, we initially collated a dataset comprising 10,000 time steps to pre-train the reward model. This model then underwent 100 pre-training iterations, a step deemed essential to adequately prepare the reward model before embarking on the principal policy learning phase. Following this initial warm-up period, the reward model was trained for 10 iterations after each new trajectory was incorporated. Moreover, to systematically gauge performance and progress, evaluations were carried out at intervals of every 5,000 time steps. The computational resources for these procedures were NVIDIA GeForce RTX 2080 Ti GPU clusters with 8GB of memory, dedicated to training and evaluating tasks.
Table 3: Performance comparison across reward bag with various-length configurations over 3 trials. In this table, “Short” refers to bags with lengths varying from 25 to 200, and “Long” denotes the setting with bag lengths from 100 to 500. The best and comparable methods based on the paired t-test at the significance level $5\%$ are highlighted in boldface.
Bag Setting Environment SAC IRCR RRD LIRPG HC RBT(ours)
Short Ant-v2 3.21 269.12 2661.56 -1407.15 20.03 5359.85
(1.97) (224.66) (1675.65) (504.59) (46.35) (129.28)
Hopper-v2 286.53 3275.05 2508.76 287.38 474.21 3433.06
(72.51) (44.68) (690.07) (114.65) (66.93) (96.89)
HalfCheetah-v2 5.92 10480.33 10382.80 1225.35 3982.43 11073.88
(23.06) (202.27) (516.85) (162.88) (433.75) (181.43)
Walker2d-v2 222.21 3840.30 3999.14 328.97 348.36 5198.09
(66.72) (666.24) (561.49) (107.69) (174.28) (225.50)
Long Ant-v2 -115.10 215.38 2600.64 -2552.99 -0.53 4897.50
(138.93) (92.94) (1229.27) (419.44) (5.07) (292.93)
Hopper-v2 360.36 3015.96 3089.52 325.84 652.28 3447.64
(118.06) (408.08) (433.23) (50.25) (92.41) (83.02)
HalfCheetah-v2 -115.56 5944.17 8591.38 -571.35 4563.99 10880.76
(35.01) (3421.46) (1048.08) (374.64) (568.19) (441.65)
Walker2d-v2 251.11 3397.93 4221.37 284.27 497.27 4979.07
(144.24) (682.19) (282.75) (7.54) (134.06) (166.95)
Appendix D Additional Experimental Results
This section provides further analysis and insights through additional experiments to complement the main findings presented in the paper.
Figure 6: Performance comparison adding SAC with reward shaping (SAC-shaping) in the Hopper-v2 and Walker2d-v2 environments, displaying both the mean and standard deviation over 6 trials with different random seeds. The experiments are conducted across six specific fixed bag lengths: 5, 25, 50, 100, 200, and 500, as well as with trajectory feedback, labeled as $9$, in each environment.
Figure 7: Ablation study of reward model components across various environments. The chart presents mean and standard deviation of rewards over 6 trials with 1e6 timesteps, showcasing the efficacy of the full proposed method relative to its variants without certain features.
Figure 8: Rewards comparison and agent states in a trajectory with a bag length of 100 in the Hopper-v2 environment. The top graph compares predicted rewards against true rewards and aggregated bagged rewards.
D.1 Experimental Result of Various-Length Reward Bags
In Table 3, we present the experiment results on various-length reward bags. The experiment depict in the table showcases the results of various methods applied across different environments with varying bag lengths of rewards, where bags are one next to another as in the definition of RLBR. This experiment reveals that longer bags tend to degrade the performance of most methods. However, our RBT method appears to be less sensitive to changes in bag length, maintaining robust performance even when the bag length is equal to the full trajectory. This result aligns with the result in Section 5.1.
D.2 Ablation Study
We conducted comparisons to examine the role of RBT’s modules. As shown in Fig. 7, the full RBT model consistently outperforms its variants, indicating a synergistic effect when all components are used together. Performance drops significantly when the bidirectional attention mechanism is removed, especially in complex environments like Ant-v2 and HalfCheetah-v2, suggesting its critical role in accurate reward prediction. Additionally, we can observe that removing the next state prediction component weakens the reward model’s understanding of environmental dynamics, reducing reward prediction accuracy and hindering policy learning. The greatest performance decline occurs when both the next state prediction and bidirectional self-attention mechanism are absent, underscoring their individual and combined importance in building a robust reward model.
D.3 Rewards Comparison in Hopper-v2 with Bagged Rewards
Fig. 8 shows a comparison of predicted rewards, true rewards, and aggregated bagged rewards for a trajectory with a bag length of 100 in the Hopper-v2 environment. It shows how well the predicted rewards align with the true and bagged rewards over the course of the trajectory, highlighting the effectiveness of the reward model.
Beneath the figure, a series of images depicts a complete jump cycle by the agent, illustrating its motion sequence: mid-air, landing, jumping, and returning to mid-air. Red boxes highlight specific states that correspond to reward peaks and troughs, representing moments of maximum, minimum, and moderate rewards. In the Hopper-v2 environment, rewards consist of a constant “healthy reward” for operational integrity, a “forward reward” for progress in the positive x-direction, and a “control cost” for penalizing large actions. At peak reward instances, the agent is typically fully grounded in an optimal posture for forward leaping, which maximizes the “forward reward” through pronounced x-direction movement. Concurrently, it sustains the “healthy reward” and minimizes “control cost” through measured, efficient actions. This analysis underscores that the RBT can adeptly decode the environmental dynamics and the nuanced reward redistribution even under the setting of RLBR.
D.4 Reward Shaping
To complement our findings, we conducted additional experiments on reward shaping, following the naive shaping approach described in Hu et al. (2020). These experiments were performed in the Hopper-v2 and Walker2d-v2 environments due to the availability of task-specific weights from the original paper. As shown in the Fig. 6, the performance of SAC with reward shaping (SAC-shaping) varies with different bag lengths.
At shorter bag lengths, SAC-shaping tends to perform worse than standard SAC, possibly because the more frequent but less informative rewards add noise to the training process. However, as bag length increases, SAC-shaping slightly outperforms SAC, likely because the shaped rewards offer clearer, longer-term signals that assist in learning effective policies when rewards are less frequent. Essentially, reward shaping at longer bag lengths creates a more stable and beneficial learning signal for policy improvement. The figure demonstrates that while reward shaping improves SAC performance with larger bag sizes, our RBT method consistently outperforms both. RBT excels because it not only smooths the rewards but also redistributes them by considering the context within each bag. This comprehensive approach allows RBT to more accurately capture environmental complexities, leading to better performance across different bag lengths, indicating its robustness and efficiency in utilizing extended sequences of data.
D.5 Architecture Sensitivity
Fig. 9 illustrates the sensitivity of our architecture to different input sequence lengths during training (Seq len) and prediction lengths during reward relabeling (Relabel len) in the Ant-v2 environment. The learning curves, based on three independent runs with random initialization, show how varying these parameters affects the performance of the agent.
Although the configuration highlighted by the red box (Seq len of 500 and Relabel len of 100) demonstrates the best performance, the results also show that our proposed model is capable of learning effectively across various other configurations. This analysis underscores the importance of tuning input sequence length during training and prediction length during reward relabeling for optimal performance, while also demonstrating the ability of the RBT model to learn under different parameter settings, showcasing its flexibility and effectiveness in reinforcement learning tasks.
Figure 9: Learning curves on Ant-v2 with different length of input sequence in training (Seq len) and predict length during relabeling process (Relabel len), based on 3 independent runs with random initialization. Within each of the smaller graphs, the curves represent results from experiments with different bag lengths. Specifically, there are three bag lengths evaluated: 25, 100, and what is labeled as $9$, which we interpret as a proxy for trajectory feedback. The graph highlighted by the red box indicates our chosen parameter setting for the experiment, which is a input sequence length of 100 and a predict length of 500.

### v3

@@
 In Reinforcement Learning (RL), it is commonly assumed that an immediate reward signal is generated for each action taken by the agent, helping the agent maximize cumulative rewards to obtain the optimal policy.
-However, in many real-world scenarios, immediate reward signals are not obtainable; instead, agents receive a single reward that is contingent upon a partial sequence or a complete trajectory.
-In this work, we define this challenging problem as Reinforcement Learning from Bagged Reward (RLBR), where sequences of data are treated as bags with non-Markovian bagged rewards.
-We provide a theoretical study to establish the connection between RLBR and standard RL in Markov Decision Processes (MDPs).
-To effectively explore the reward distributions within these bags and enhance policy training, we propose a Transformer-based reward model, the Reward Bag Transformer, which employs a bidirectional attention mechanism to interpret contextual nuances and temporal dependencies within each bag.
-Our empirical evaluations reveal that the challenge intensifies as the bag length increases, leading to the performance degradation due to reduced informational granularity.
-Nevertheless, our approach consistently outperforms existing methods, demonstrating the least decline in efficacy across varying bag lengths and excelling in approximating the original MDP’s reward distribution.
+However, in many real-world scenarios, designing immediate reward signals is difficult; instead, agents receive a single reward that is contingent upon a partial sequence or a complete trajectory.
+In this work, we define this challenging problem as RL from Bagged Reward (RLBR), where sequences of data are treated as bags with non-Markovian bagged rewards, leading to the formulation of Bagged Reward Markov Decision Processes (BRMDPs).
+Theoretically, we demonstrate that RLBR can be addressed by solving a standard MDP with properly redistributed bagged rewards allocated to each instance within a bag.
+Empirically, we find that reward redistribution becomes more challenging as the bag length increases, due to reduced informational granularity.
+Existing reward redistribution methods are insufficient to address these challenges.
+Therefore, we propose a novel reward redistribution method equipped with a bidirectional attention mechanism, enabling the accurate interpretation of contextual nuances and temporal dependencies within each bag.
+We experimentally demonstrate that our proposed method consistently outperforms existing approaches.
 The code is available at an anonymous link: https://anonymous.4open.science/r/RLBR-F66E/.
@@
 One common and essential assumption for most RL algorithms is the availability of immediate reward feedback at each time step of the decision-making process.
-However, this assumption is violated in many real-world applications.
+However, designing such immediate reward feedback is quite difficult and may mislead the policy learning in many real-world applications Kwon et al.
+(2023); Lee et al.
+(2023).
 Recognizing this gap, numerous studies Watkins (1989); Ke et al.
@@
 (2021); Zhang et al.
-(2023b) have explored the concept of delayed rewards, primarily focusing on trajectory feedback where rewards are allocated at the end of a sequence.
-Similarly, sparse reward settings, where agents receive infrequent and instance-specific feedback, present significant challenges in the well-known exploration-exploitation trade-off Riedmiller et al.
-(2018); Devidze et al.
-(2022); Rengarajan et al.
-(2022).
-On the other hand, real-world applications, such as autonomous driving (see Fig.
-1), often feature complex non-immediate reward structures that neither sparse rewards nor trajectory feedback can fully capture.
-Providing reward for every action is impractical, and focusing only on end goals ignores crucial aspects of the journey.
+(2023) have explored the concept of RL with Trajectory Feedback (RLTF), primarily focusing on trajectory feedback where rewards are allocated at the end of a sequence.
+However, real-world applications, such as autonomous driving (see Fig.
+1), often feature complex non-immediate reward structures that RLTF cannot fully capture.
+Providing rewards for every action is impractical, while focusing only on end goals ignores crucial aspects of the journey.
 Typically, rewards are linked to completing specific tasks or sequences of actions, rather than individual actions or the final objective Early et al.
 (2022); Gaon and Brafman (2020).
-However, previous studies (see Fig.
-2) mainly focused on learning desirable policies with immediate rewards Sutton and Barto (1998); Vinyals et al.
-(2019), trajectory feedback Watkins (1989); Arjona-Medina et al.
-(2019), or sparse rewards where feedback is infrequent and tied only to the current instance Riedmiller et al.
-(2018); Devidze et al.
-(2022); Rengarajan et al.
-(2022), tending to fail under such scenarios.
-To address these challenges, we introduce Reinforcement Learning from Bagged Rewards (RLBR), which better aligns with real-world scenarios by considering the cumulative effect of a series of state-action pairs.
-In RLBR, we define sequences of state-action pairs as bags, each associated with a cumulative bagged reward.
-This framework includes both the traditional RL setting, where each bag only contains a single instance (the first line of Fig.
-2), and the trajectory feedback setting, where a bag spans an entire trajectory (the third line of Fig.
-2).
+Previous studies focused on learning desirable policies with immediate rewards (Sutton and Barto, 1998; Vinyals et al., 2019) or trajectory feedback (Arjona-Medina et al., 2019) are not ideal methods for such scenarios.
+To address these challenges, we introduce RL from Bagged Rewards (RLBR), a framework that better aligns with real-world scenarios by considering the cumulative effect of a series of instances.
+In RLBR, sequences of instances are defined as bags, each associated with a bagged reward.
+This framework encompasses both the traditional RL setting, where each bag contains only a single instance, and the trajectory feedback setting, where a bag spans the entire trajectory, as special cases.
 Furthermore, RLBR offers the potential to reduce the labeling workload by lessening the frequency of reward annotations.
 However, this benefit is balanced by increased learning complexity due to the reduced granularity of information.
-Figure 1: An example of the reward bag structure on an autonomous driving trajectory.
-Each segment of the driving sequence is evaluated and assigned a score by an evaluator, and the score for each sequence is integrated based on the performance at each step.
+Figure 1: An example of the reward bag structure on an autonomous driving trajectory. Each segment of the driving sequence is evaluated and assigned a score by an evaluator.
 In the RLBR framework, our focus is on leveraging bagged reward information to discern the significance of each instance within a bag and to understand the relationships among different bags.
@@
 Our research contributes to the field in several ways.
-First, we establish the RLBR framework as a general problem setting and connect it theoretically to traditional Markov Decision Processes (MDPs), elaborated in Section 3.
-In Section 4, we introduce a Transformer-based reward model designed to assign rewards to individual instances while capturing environmental dynamics.
-Additionally, we propose an algorithm that alternates between optimizing this reward model and the policy, thereby enhancing the effectiveness of both components.
+First, we establish the RLBR framework as a general problem setting by formulating it with an extension of the traditional Markov Decision Processes (MDPs), the Bagged Reward MDPs (BRMDPs), as elaborated in Section 3.
+In Section 4, we theoretically connect MDPs after reward redistribution to the original BRMDPs, which motivates us to employ a bidirectional attention mechanism within a Transformer-based framework to assign rewards to individual instances while capturing environmental dynamics.
+Additionally, we propose an algorithm that alternates between optimizing this reward redistributing model and the policy, thereby enhancing the effectiveness of both components.
 Finally, in Section 5, our experiments show that the performance of baseline methods drops as the bag length increases, indicating that the larger bag length will increase the learning difficulty.
-Furthermore, we experimentally demonstrate the superiority of our method through comparative performance analyses and validate the ability of the proposed model to mimic the reward distribution of the ground truth MDP, highlighting its contextual understanding and adaptability to environmental dynamics.
+Furthermore, we experimentally demonstrate the superiority of the proposed method through comparative performance analyses and validate the ability of the proposed model to mimic the reward distribution of the ground truth MDP, highlighting its contextual understanding and adaptability to environmental dynamics.
 2 Related Work
-RL with Trajectory Feedback.
-RL with Trajectory Feedback (RLTF), termed episodic rewards or delayed rewards in some works (see the third line in Fig.
-2), has become increasingly prominent in many applications (Watkins, 1989; Arjona-Medina et al., 2019; Zhang et al., 2023b).
-A key approach to this challenge is reward redistribution, aiming to assign rewards to individual instances more effectively.
+With the growing attention on RLTF, where trajectory feedback is often considered the sum of rewards over the entire trajectory, new methods specifically designed for RLTF have emerged.
 Return Decomposition for Delayed Rewards (RUDDER) Arjona-Medina et al.
-(2019) used a return-equivalent formulation for precise credit assignment, with advancements incorporating expert demonstrations and language models Patil et al.
-(2022); Liu et al.
+(2019) used a return-equivalent formulation for precise credit assignment, with advancements incorporating expert demonstrations and language models Liu et al.
 (2019); Widrich et al.
+(2021); Patil et al.
 (2022).
-Iterative Relative Credit Refinement (IRCR) Gangwani et al.
-(2020) presented a uniform reward redistribution model based on equal contributions from all state-action pairs, while Randomized Return Decomposition (RRD) Ren et al.
-(2021) proposed a novel upper bound for return-equivalent assumptions, integrating return decomposition with uniform reward redistribution.
+Both Iterative Relative Credit Refinement (IRCR) Gangwani et al.
+(2020) and Randomized Return Decomposition (RRD) Ren et al.
+(2021) assumed that each state-action pair contributes equally to the trajectory feedback.
+Specifically, IRCR presented a uniform reward redistribution model, while RRD proposed a novel upper bound for return-equivalent assumptions, integrating return decomposition with uniform reward redistribution.
 Additionally, Han et al.
 (2022) modified RL algorithms to use sequence-level information, helping agents learn from broader structures and long-term outcomes.
-However, our focus is on the original, unobservable rewards within reward bags, not the aggregated bagged rewards.
-Our experiments (see Section 5) show that the method by Han et al.
-(2022) is ineffective for long sequences.
-Following the previous methodology, we adopt a reward redistribution learning strategy to enhance policy learning in the context of reward bags.
-However, conventional reward redistribution methods fail to effectively extract information from bag-level rewards, as they did not take the structure of the bag into consideration.
-Our experimental results demonstrate that directly applying previous methods to the reward bag setting does not yield results as promising as those achieved by our algorithm.
-Figure 2: Comparison of four different types of reward settings in RL.
-In traditional RL, a reward is given based on each instance Sutton and Barto (1998); Vinyals et al.
-(2019), whereas in sparse reward settings, only some instances receive rewards Riedmiller et al.
-(2018); Devidze et al.
-(2022); Rengarajan et al.
+On the other hand, some studies on sparse rewards, such as reward shaping Ng et al.
+(1999); Tambwekar et al.
+(2019); Hu et al.
+(2020) and intrinsic reward methods Sorg et al.
+(2010); Pathak et al.
+(2017); Zheng et al.
+(2018, 2021), could be directly applied to RLTF.
+However, these approaches were primarily focused on addressing the exploration-exploitation trade-off in sparse reward settings, which made them less ideal for solving RLTF problems.
+A detailed discussion on the sparse reward setting and methods is provided in “Discussion on Sparse Rewards” in Appendix.
+However, previous RLTF works that assume each state-action pair contributes equally to the trajectory feedback did not consider that instances contributing more to the feedback should receive greater attention.
+On the other hand, methods that consider sequence information struggled to capture effective information from long trajectories, resulting in suboptimal training outcomes, as evidenced by prior studies Ren et al.
+(2021); Zhang et al.
 (2023).
-Both types of rewards are Markovian.
-Trajectory feedback and the proposed bagged reward both address non-Markovian situations; however, trajectory feedback provides only one reward signal for the entire trajectory Watkins (1989); Arjona-Medina et al.
-(2019), while bagged rewards can include multiple reward signals within the sequence.
-RL with Sparse Rewards.
-The sparse reward setting (see the second line in Fig.
-2) presents significant challenges due to infrequent feedback, making it difficult for agents to effectively explore the environment and discover successful strategies.
-To address this challenge, various methods have been developed to enhance exploration.
-Reward shaping strategies Ng et al.
-(1999); Hu et al.
-(2020); Tambwekar et al.
-(2019) added rewards to actions in a way that guides the agent towards better policies without altering the original reward function.
-Curiosity-driven methods Pathak et al.
-(2017); Sekar et al.
-(2020) encouraged agents to explore the environment by visiting unseen states, potentially solving tasks with sparse rewards.
-Additionally, curriculum learning in RL Florensa et al.
-(2018); Riedmiller et al.
-(2018) involved presenting an agent with a sequence of tasks with gradually increasing complexity, allowing the agent to eventually solve the initially given sparse reward task.
-In contrast to sparse rewards, which are typically infrequent Markovian signals associated with individual state-action pairs, bagged rewards depend on the cumulative effect of sequences of state-action pairs and are non-Markovian, requiring context from the entire sequence to accurately assess the contribution of each instance.
-Transformers for RL.
-Transformers, introduced by Vaswani et al.
-(2017), have proven effective in RL environments requiring high sample efficiency and generalization, such as StarCraft Vinyals et al.
-(2019); Zambaldi et al.
-(2019) and DMLab30 Parisotto et al.
-(2020).
-They have also been used for sequential modeling in offline RL Chen et al.
-(2021); Janner et al.
-(2021b) and as reward models in policy learning from offline datasets Kim et al.
-(2023).
-Luo et al.
-(2021) combined deep convolution transfer learning models and inverse RL for reward function acquisition, while Zhang et al.
-(2023a) transformed non-Markovian reward processes into Markovian ones, enhancing online interaction efficiency.
-For reward bags, where rewards are linked to sequences of state-action pairs, Transformers are ideal due to their ability to handle long-range dependencies.
-Our approach leverages a Transformer-based model to redistribute bag-level rewards into instance-level rewards, capitalizing on the Transformer’s proficiency in sequential data analysis and attention mechanisms.
+Building on previous methods, we adopt a reward redistribution learning strategy to enhance policy learning in the context of reward bags.
+The proposed method aims to extract information from bag-level rewards, taking into account the contribution of each instance, ultimately learning an effective policy.
+Experimental results (see Section 5) demonstrate that the proposed method can effectively capture information from long sequences, and the rich information from the bags makes it more effective.
 3 Reinforcement Learning from Bagged Reward
-In this section, we first provide preliminaries on RL with immediate rewards and trajectory feedback, and then formulate the RLBR problem with an extension of the traditional MDP, the Bagged Reward MDP (BRMDP).
-We further conduct a theoretical analysis of the relationship between the BRMDP and the traditional MDP, showing that solving RLBR problems not only relaxes the demands on reward acquisition but also guarantees finding the optimal policy even with limited reward information.
+In this section, we first provide preliminaries on RL with immediate rewards and trajectory feedback, and then formulate the RLBR problem with an extension of the traditional MDPs, the BRMDPs.
 3.1 Preliminaries
-We consider finite-horizon RL in this paper, which is traditionally modeled using a finite MDP, where rewards are promptly provided for each state-action pair (Sutton and Barto, 1998).
+In traditional RL problems, which are modeled using finite MDPs, each state-action pair is promptly associated with a reward (Sutton and Barto, 1998).
 This paradigm is encapsulated in a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathit{P},\mathit{r},\mu)$, with $s\in\mathcal{S}$ and $a\in\mathcal{A}$ as the sets of states and actions, $\mathit{P}$ as the state transition probability function, $\mathit{r}$ as the immediate reward function, and $\mu$ as the initial state distribution.
@@
 $J(\pi)=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\bigg{]}.$ (1)
-Distinct from this traditional approach, RLTF offers feedback only after a complete sequence of actions, or a trajectory (Arjona-Medina et al., 2019; Zhang et al., 2023b).
-A trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T-1},a_{T-1})\}$ includes $T$ state-action pairs, with a reward $R_{\mathrm{traj}}(\tau)$ that is the sum of latent immediate rewards $\sum_{t=0}^{T-1}r(s_{t},a_{t})$, but only the cumulative reward is observable at the end of the trajectory.
+Given that associating each state-action pair with a reward is challenging in real-world scenarios and involves significant labeling costs, RLTF, also known as episodic or delayed rewards in some works, has become increasingly prominent in many applications (Watkins, 1989; Arjona-Medina et al., 2019; Zhang et al., 2023).
+Distinct from traditional RL, RLTF offers only one feedback after a complete trajectory (Arjona-Medina et al., 2019; Zhang et al., 2023).
+A trajectory $\tau=\{(s_{0},a_{0}),(s_{1},a_{1}),\ldots,(s_{T-1},a_{T-1})\}$ consists of $T$ state-action pairs, with a cumulative reward $R_{\mathrm{traj}}(\tau)$ that is the sum of latent immediate rewards $\sum_{t=0}^{T-1}r(s_{t},a_{t})$, observable only at the end.
 Denoting by $\mathcal{T}(\pi)$ as the distribution of trajectories induced by $\pi,P,\mu$, the learning objective in RLTF is to maximize the expected trajectory-based reward:
-$J_{\mathrm{traj}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\big{[}R_{\mathrm{ep}}(\tau)\big{]}.$ (2)
+$J_{\mathrm{traj}}(\pi)=\mathbb{E}_{\mathcal{T}(\pi)}\big{[}R_{\mathrm{traj}}(\tau)\big{]}.$ (2)
 3.2 Problem Formulation
-We formally define the setting of RLBR, which has a granularity of rewards in between the above two settings.
+We formulate the RLBR problem as an extension of the traditional MDP, the BRMDP, which has a granularity of rewards in between the above two settings.
 First, we define the notion of bags, which are sub-pieces of complete trajectories.
@@
   • $\mathit{P}$ is the state transition probability function.
-  • $\mathit{R}$ denotes the reward function over bagged reward: $\mathit{R}(B_{i,n_{i}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})$.
+  • $\mathit{R}$ denotes the bagged reward function, which define the reward over a bag: $\mathit{R}(B_{i,n_{i}})$.
   • $\mu$ represents the initial state distribution.
-The essential properties of BRMDP are two-fold: 1) only bagged rewards can be obtained by the learner, not immediate rewards; 2) the bagged reward is the accumulated sum of immediate rewards within the bag.
-In the RLBR framework, a bag $B_{i,n_{i}}$ metaphorically aggregates individual rewards from a contiguous sequence of state-action pairs into a unified reward unit.
+In the RLBR framework, a bag $B_{i,n_{i}}$ metaphorically a unified reward unit for a contiguous sequence of state-action pairs.
 A trajectory $\tau$ is a composite of a set of bags, denoted as $\mathcal{B}_{\tau}$, which ensures that each trajectory includes at least one reward bag.
@@
 This adaptability highlights the capacity of the RLBR framework to accommodate varying reward structure scenarios.
-3.3 Equivalence of Optimal Policies in the BRMDP and the Original MDP
-In the BRMDP, agents are unable to directly observe the immediate reward linked to each state-action pair.
-Nonetheless, we aim for agents to optimize policies in environments where bagged rewards are accessible and to perform well on the original task.
-To unify the learning objectives of the BRMDP with those of the traditional MDP, we formulate a theorem that crystallizes this relationship.
-The significance of this theorem lies in its assertion that the optimal policy derived in a BRMDP setting is congruent with that of an MDP when the bagged rewards are considered as cumulative sums.
-This finding not only validates the theoretical underpinnings of our approach but also affirms its practical relevance across various RL paradigms.
-Theorem 1.
-Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag.
-In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
-The detailed proof of Theorem 1 is deferred to Appendix B.
-By this theorem, we can see that optimizing a policy on a BRMDP is equivalent to optimizing on the original MDP.
-The crux of the matter is how to accurately redistribute the bagged rewards so that the policies learned under these distributed rewards are more close to the policies on the original MDP.
 Figure 2: The illustration of the Reward Bag Transformer (RBT) architecture.
@@
 The bidirectional attention layer is used for reward redistribution, utilizing the outputs of the Causal Transformer to predict instance-level rewards.
-Predicting the next state helps the model understand the environment, thereby improving reward prediction.
-4 Reward Bag Transformer
-This section delves into the intricacies of leveraging the bidirectional attention mechanism Seo et al.
-(2016); Vaswani et al.
-(2017); Devlin et al.
-(2019) for reward redistribution, designs a Transformer architecture, and outlines a comprehensive algorithm for the cultivation of efficient policies under the BRMDP framework.
+The next state prediction helps the model understand the environment, thereby improving reward prediction.
+4 Reward Redistribution for RLBR
+This section delves into the proposed reward redistribution method for RLBR.
+We aim for the reward redistribution to meet the following criteria: first, the policies learned from the MDP after reward redistribution should be consistent with those learned from the BRMDP; second, the redistribution process should reflect each instance’s contribution based on the sequence information.
+The subsequent subsections are structured to first theoretically prove the rationality of reward redistribution, followed by proposing the use of a bidirectional attention mechanism for this redistribution, and finally outline a comprehensive algorithm for developing efficient policies within the BRMDP framework.
+4.1 Equivalence of Optimal Policies in the BRMDP and the MDP after Reward Redistribution
+In the BRMDP, agents do not have direct access to the immediate rewards associated with each instance.
+To address this, we adopt a common approach used in trajectory feedback scenarios, modeling the cumulative sum form as the reward redistribution strategy to convert bagged rewards into instance-level rewards Ren et al.
+(2021); Arjona-Medina et al.
+(2019); Zhang et al.
+(2023).
+Specifically, we define $\mathit{R}(B_{i,n_{i}})=\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t})$, where $\hat{r}(s_{t},a_{t})$ represents the redistributed reward.
+However, a critical question arises: How can we ensure that the policy learned with redistributed rewards performs well in the original BRMDP?
+To address this, we provide the following theorem:
+Theorem 1.
+Provide a BRMDP where the bagged reward is regarded as the sum of rewards for state-action pairs of a MDP after reward redistribution contained within the same bag.
+In this context, the set of optimal policies for the BRMDP $\Pi_{\mathrm{B}}$ aligns with that of the MDP after reward redistribution $\Pi$, implying that $\Pi=\Pi_{\mathrm{B}}$.
+The detailed proof of Theorem 1 is deferred to the “Omitted Proof” in Appendix.
+By this theorem, we can see that optimizing a policy on a MDP after reward redistribution is equivalent to optimizing on the original BRMDP.
 4.2 Reward Redistribution based on Bidirectional Attention Mechanism
-Building on the foundation of Theorem 1 that optimal policies in the BRMDP and MDP are equivalent, our focus shifts to the crucial process of redistributing rewards within a bag from the BRMDP.
+Building on Theorem 1 that optimal policies in the BRMDP and the MDP after reward redistribution are equivalent, our focus shifts to the crucial process of redistributing rewards within a bag.
 To capture the contextual influence of each instance within the sequence, the Causal Transformer Vaswani et al.
@@
 (2021); Janner et al.
-(2021a); Micheli et al.
+(2021); Micheli et al.
 (2023), where only previous instances influence the current prediction due to the unobservability of future instances.
-However, given that our bagged rewards are non-Markovian and both preceding and subsequent instances affect the contribution of the current instance to the bagged reward, a bidirectional attention mechanism Seo et al.
+However, since our bagged rewards are generally non-Markovian, both preceding and subsequent instances influence the contribution of the current instance to the bagged reward, understanding the relationships among instances within a bag is therefore pivotal.
+This insight leads us to employ a bidirectional attention mechanism Seo et al.
 (2016); Vaswani et al.
 (2017); Devlin et al.
-(2019) becomes pivotal.
+(2019) to capture these relationships.
 This mechanism connects both past and future instances within a bag, enabling a more comprehensive understanding of contextual influences.
-By quantitatively evaluating the contribution of each instance, the bidirectional attention mechanism facilitates nuanced and effective reward redistribution.
-4.2 Proposed Approach
-We introduce the Reward Bag Transformer (RBT), a novel approach designed for the BRMDP framework.
+By quantitatively evaluating the contribution of each instance, the bidirectional attention mechanism facilitates nuanced and effective reward redistribution, as experimentally demonstrated in Section 5.
+4.3 Reward Bag Transformer
+We introduce the Reward Bag Transformer (RBT), a novel reward redistribution method for the BRMDP.
 The RBT is engineered to comprehend the complex dynamics of the environment through bags and to precisely predict instance-level rewards, facilitating effective reward redistribution.
-Causal Transformer for Reward Representation.
+4.3.1 Causal Transformer for Reward Representation
 Referring to Fig.
-3, the RBT comprises a Causal Transformer (Vaswani et al., 2017; Radford et al., 2018), which maintains the chronological order of state-action pairs Chen et al.
+2, the RBT comprises a Causal Transformer (Vaswani et al., 2017; Radford et al., 2018), which maintains the chronological order of state-action pairs Chen et al.
 (2021); Janner et al.
-(2021a).
+(2021).
 For each time step $t$ in a sequence of $M$ time steps, the Causal Transformer, represented as a function $f$, processes the input sequence $\sigma=\{s_{0},a_{0},\ldots,s_{M-1},a_{M-1}\}$, generating the output $\{x_{t}\}_{t=0}^{M-1}=f(\sigma)$.
-By aligning the output head $x_{t}$ with the action token $a_{t}$, we directly model the consequence of actions, which are pivotal in computing immediate rewards and predicting subsequent states, which in turn aids the model in better understanding the environmental dynamics.
-Bidirectional Attention Layer for Reward Redistribution.
+By aligning the output head $x_{t}$ with the action token $a_{t}$, we directly model the consequence of actions, which are pivotal in computing immediate rewards and predicting subsequent states, thereby helping the model better understand environmental dynamics.
+4.3.2 Bidirectional Attention Layer for Reward Redistribution
 Once we have obtained the output embeddings $\{x_{t}\}_{t=0}^{M-1}$, for reward prediction, they pass through a bidirectional attention layer to produce $\{\hat{r}_{t}\}_{t=0}^{M-1}$, where $\hat{r}_{t}\equiv\hat{r}_{\theta}(s_{t},a_{t})$ with $\theta$ being the RBT parameters.
@@
 The core of the RBT architecture is its bidirectional attention mechanism.
-For each output embedding $x_{t}$, we apply three different linear transformations to obtain embeddings for query $\mathbf{q}_{t}\in\mathbb{R}^{d}$, key $\mathbf{k}_{t}\in\mathbb{R}^{d}$, and value $v_{t}\in\mathbb{R}$.
+For each output embedding $x_{t}$, we apply three different linear transformations to obtain embeddings for query $\mathbf{q}_{t}\in\mathbb{R}^{d}$, key $\mathbf{k}_{t}\in\mathbb{R}^{d}$, and value $v_{t}\in\mathbb{R}$, where $d$ is the embedding dimension of the key.
 Then the instance-level reward is calculated by
-$\hat{r}_{t}=\sum_{\ell=0}^{M-1}\texttt{softmax}(\frac{\{\langle\mathbf{q}_{t},\mathbf{k}_{t^{\prime}}\rangle\}_{t^{\prime}=0}^{M-1}}{\sqrt{d}})_{\ell}\cdot v_{\ell},$ (4)
-where $d$ is the embedding dimension of the key.
+$\hat{r}_{t}=\sum_{\ell=0}^{M-1}\texttt{softmax}(\frac{\{\langle\mathbf{q}_{t},\mathbf{k}_{t^{\prime}}\rangle\}_{t^{\prime}=0}^{M-1}}{\sqrt{d}})_{\ell}\cdot v_{\ell}.$ (4)
 The rescaling operation is used to prevent extremely small gradients as in Vaswani et al.
 (2017).
-This mechanism enables the RBT to consider both the immediate and contextual relevance of each state-action pair in the trajectory when predicting rewards.
+This mechanism enables the RBT to consider both the immediate and contextual relevance of each instance in a bag when predicting rewards.
 Algorithm 1 Policy Optimization with RBT
@@
 3:     Store trajectory $\tau$ with bag information $\{(B_{i,n_{i}},R(B_{i,n_{i}}))\}_{B_{i,n_{i}}\in\mathcal{B}_{\tau}}$ in $\mathcal{D}$.
-4:     Sample batches from $\mathcal{D}$ and estimate bag loss based on Eq. (7).
-5:     Update RBT parameters $\theta$ based on the estimated loss.
-6:     Relabel rewards in $\mathcal{D}$ using the updated RBT.
-7:     Optimize policy using the relabeled data by off-the-shelf RL algorithms (such as SAC).
-8:  end for
+4:     Sample batches from $\mathcal{D}$.
+5:     Estimate bag loss based on Eq. (7).
+6:     Update RBT parameters $\theta$ based on the estimated loss.
+7:     Relabel rewards in $\mathcal{D}$ using the updated RBT.
+8:     Optimize policy using the relabeled data by off-the-shelf RL algorithms (such as SAC Haarnoja et al. (2018)).
+9:  end for
 4.4 Learning Objectives
@@
 These objectives are critical for enabling the model to navigate the complex dynamics of BRMDP environments.
-Reward Prediction.
+4.4.1 Reward Prediction
 The RBT is trained to ensure that, for each reward bag, the sum of predicted instance-level rewards matches the total bagged reward.
 This is vital for maintaining the integrity of the reward structure in the BRMDP framework.
-The loss function for this objective is expressed as
-$\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})=\left(\sum_{t=i}^{i+n_{i}-1}\hat{r}_{t}-R(B_{i,n_{i}})\right)^{2},$ (5)
-where $B_{i,n_{i}}$ represents the reward bag starting at time step $i$ with length $n_{i}$, and $R(B_{i,n_{i}})$ is the total reward for this bag.
+The loss function is expressed as
+$\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})=\left(\sum_{t=i}^{i+n_{i}-1}\hat{r}_{t}-R(B_{i,n_{i}})\right)^{2}.$ (5)
 This formulation encourages the RBT to learn a nuanced distribution of rewards across states and actions within a bag.
 At the same time, it ensures that the sum of redistributed rewards matches the total bagged reward, maintaining consistency as per Theorem 1.
-State Transition Forecasting.
+4.4.2 State Transition Forecasting
 Alongside reward prediction, the RBT is tasked with accurately predicting the next state in the environment given the current state and action.
@@
 […]
-Composite Loss. The final learning objective combines the reward and state prediction losses:
+4.4.3 Composite Loss
+The final learning objective combines the reward and state prediction losses:
 $\mathcal{L}_{\mathrm{bag}}$ $(\theta)=\underset{\tau\sim\mathcal{D}}{\mathbb{E}}\big{[}\mathcal{L}_{\mathrm{r}}(\theta;B_{i,n_{i}})\big{|}B_{i,n_{i}}\in\mathcal{B}_{\tau}\big{]}$ (7)
@@
 The RBT’s dual predictive capacity is its key advantage, enabling precise reward redistribution to individual instances and forecasting the next state.
-This leverages environmental dynamics for enhanced reward distribution as experimentally evidenced in the ablation study in Appendix D.2.
+This leverages environmental dynamics for enhanced reward distribution as experimentally evidenced in Section 5.
 Integrated with off-the-shelf RL algorithms such as SAC Haarnoja et al.
@@
 5 Experiment
-In the following experimental section, we scrutinize the efficacy of our proposed method using benchmark tasks from both the MuJoCo Brockman et al.
+In the following experiment section, we scrutinize the efficacy of our proposed method using benchmark tasks from both the MuJoCo Brockman et al.
 (2016) and the DeepMind Control Suite Tassa et al.
 (2018) environments, focusing on scenarios with bagged rewards.
-We assess the performance of our method to understand its overall effectiveness and examine whether the RBT reward model accurately predicts rewards.
-5.1 Compare with SOTA Methods
-Experiment Setting.
-We evaluated our method on eight benchmark tasks from both the MuJoCo locomotion suite, including Ant-v2, Hopper-v2, HalfCheetah-v2, and Walker2d-v2, and the DeepMind Control Suite, including cheetah-run, quadruped-walk, fish-upright, and cartpole-swingup.
+Initially, we assess the performance of our method to understand its overall effectiveness.
+Subsequently, we examine whether the proposed RBT reward model accurately predicts rewards.
+Finally, we evaluate the indispensability of each component of the reward model, questioning if every part is essential.
+5.1 Performance Comparison
+5.1.1 Experiment Setting
+We evaluated our method on benchmark tasks from the MuJoCo locomotion suite (Ant-v2, Hopper-v2, HalfCheetah-v2, and Walker2d-v2) and the DeepMind Control Suite (cheetah-run, quadruped-walk, fish-upright, cartpole-swingup, ball_in_cup-catch, and reacher-hard).
 Differing from standard environments where rewards are assigned at each step, our approach involved assigning a cumulative reward at the end of each bag while assigning a reward of zero to all other state-action pairs within the bag.
 The maximum length for each episode was fixed at 1000 steps across all tasks.
-Baselines. In the comparative analysis, our framework was rigorously evaluated against several leading algorithms in the domain of RL with delayed reward:
+Figure 3: Performance comparison across fixed-length reward bag in MuJoCo (top row) and DeepMind Control Suite (bottom row) environments with six different length bag settings (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$).
+The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
+5.1.2 Baselines
+In the comparative analysis, our framework was rigorously evaluated against several leading algorithms in the domain of RL with delayed reward:
   • SAC (Haarnoja et al., 2018): It directly utilized the original bagged reward information for policy training using the SAC algorithm.
-  • IRCR Gangwani et al.
-(2020): It adopted a non-parametric uniform reward redistribution approach by using the sum of immediate per-step rewards as a stand-in for trajectory returns.
-We have adapted IRCR for bagged reward setting.
+  • IRCR Gangwani et al. (2020): It adopted a non-parametric uniform reward redistribution approach. We have adapted IRCR for bagged reward setting.
   • RRD Ren et al. (2021): It employed a reward model trained with a randomized return decomposition loss. We have adapted RRD for bagged reward setting.
+  • HC Han et al.
+(2022): The HC-decomposition framework was utilized to train the policy using a value function that operates on sequences of data.
+We employed the code as provided by the original paper.
+  • Shaping Hu et al. (2020): A widely used reward shaping method for sparse reward setting, the shaping reward function is adopted from Hu et al. (2020).
   • LIRPG Zheng et al.
@@
 We use the same code provided by the paper.
-  • HC Han et al.
-(2022): The HC-decomposition framework was utilized to train the policy using a value function that operates on sequences of data.
-This approach decoupled the value function approximation task for the current step from the historical trajectory.
-We employed the code as provided by the original paper.
-While methods like RUDDER Arjona-Medina et al.
-(2019) and Align-RUDDER Patil et al.
-(2022) are known for addressing the problem of trajectory feedback, previous studies Gangwani et al.
-(2020); Ren et al.
-(2021); Zhang et al.
-(2023b) have shown superior performance using referenced methods.
-Additionally, since Align-RUDDER relies on successful trajectories for scoring state-action pairs, which is impractical in MuJoCo Patil et al.
-(2022), we ultimately excluded both methods from our comparison.
-Besides, detailed descriptions of the model parameters and hyper-parameters used during training are provided in Appendix C.
-More experimental results are included in Appendix D.1.
-Figure 4: Performance comparison in MuJoCo (top row) and DeepMind Control Suite (bottom row) environments with six different fixed-length reward bag settings (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$).
-The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
 Table 1: Performance comparison across arbitrary reward bag configurations over 6 trials with 1e6 time steps for training, presenting average scores and standard deviations.
@@
 (49.81) (937.42) (79.26) (98.05) (155.63) (248.35)
-5.1.1 Fixed-Length Reward Bags
-In the fixed-length reward bag experiment, we conducted experiments with six bag lengths (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$) across each environment.
-The experimental design followed the problem setting, which assumed that the conclusion of one reward bag directly preceded the beginning of the next.
-It aimed to illustrate the influence of varying bag lengths on the results, providing insight into how bag size affected the performance of the learning algorithm within these environments.
-As shown in Fig.
-4, the SAC method, using bagged rewards directly from the environment, suffers from a lack of guidance in agent training due to missing reward information.
+While methods like RUDDER Arjona-Medina et al.
+(2019) and Align-RUDDER Patil et al.
+(2022) are known for addressing the problem of trajectory feedback, previous studies Gangwani et al.
+(2020); Ren et al.
+(2021); Zhang et al.
+(2023) have shown superior performance using referenced methods.
+Additionally, since Align-RUDDER relies on successful trajectories for scoring state-action pairs, which is impractical in MuJoCo Patil et al.
+(2022), we ultimately excluded both methods from our comparison.
+Sparse reward baselines were included in our comparisons to highlight the distinctions and challenges unique to our bagged reward setting.
+While both settings provide rewards after a sequence, sparse rewards are Markovian, whereas our bagged rewards are non-Markovian.
+Therefore, including sparse reward baselines illustrates that directly applying sparse reward methods to bagged rewards is possible but not effective.
+For more discussion on sparse rewards, see “Discussion on Sparse Rewards” in Appendix.
+Besides, detailed descriptions of the model parameters and hyper-parameters used during training are provided in “Experiment Settings and Implementation Details”, with more experimental results available in “Additional Experimental Results” in Appendix.
+5.1.3 Evaluation Metric
+We report the average accumulative reward across 6 seeds with random initialization to demonstrate the performance of evaluated methods.
+Higher accumulative reward in evaluation indicates better performance.
+5.1.4 Experimental Results
+In the fixed-length reward bag experiment (see Fig.
+3), we conducted experiments with six bag lengths (5, 25, 50, 100, 200, and 500) and trajectory feedback (labeled as $9$) across each environment, where bagged reward is made by accumulating immediate rewards within bags.
+This aimed to illustrate the influence of varying bag lengths, providing insight into how bag size affected the performance of the learning algorithm.
+The SAC method, using bagged rewards directly from the environment, suffers from a lack of guidance in agent training due to missing reward information.
 This issue worsens with longer bag lengths, indicating that increased reward sparsity leads to less effective policy optimization.
@@
 However, notable variance in their results indicates potential consistency and reliability issues.
-The LIRPG exhibits subdued performance across tasks, as it is proposed to solve sparse reward problems (Zheng et al., 2018), which is Markovian and does not align the reward bag setting.
 The HC method excels only with shorter bag lengths, suggesting that this value function modification method struggles to utilize information from longer sequences.
+The Shaping and LIRPG methods exhibit subdued performance across tasks, as they are designed to solve sparse reward problems with Markovian rewards, which do not align with the reward bag setting.
 The proposed RBT method consistently outperforms the other approaches across all the environments and bag lengths, showing that it is not only well-suited for environments with short reward bags but also capable of handling large reward bag scenarios.
 This demonstrated the capability of RBT to learn from the sequence of instances and, by integrating bagged reward information, accurately allocate rewards to instances, thereby guiding better policy training.
-5.1.2 Arbitrary Reward Bags
-To validate the effectiveness of our approach under more complex conditions, we designed an experiment that allowed for overlaps or gaps between reward bags, and the length of each bag was no longer fixed.
+To validate the effectiveness of our approach under more complex conditions, we designed an experiment with arbitrary-length reward bags, allowing for overlaps or gaps between reward bags.
 This setup simulated more realistic scenarios and tested the robustness of our method.
-The results, as detailed in Table 1, affirm the superior performance of the proposed RBT method in these complex reward settings.
-Notably, other baseline methods that rely on reward model training experience a significant drop in effectiveness, primarily due to sample scarcity impacting model accuracy and policy learning.
-In contrast, the IRCR method, which distributes rewards uniformly without a model, maintains its efficacy in some environments.
-This outcome suggests that while approximate rewards can still guide policy learning, incorrect rewards can be highly detrimental.
-The consistent success of our approach in various reward bag scenarios indicates its potential for application to a broader range of reward structures, highlighting its versatility and robustness in handling more intricate reward dynamics.
+The results, detailed in Table 1, confirm the superior performance of the proposed RBT method in these complex reward settings.
+These findings suggest RBT’s potential for broader applications, demonstrating its versatility and robustness in handling intricate reward dynamics.
+We also conducted experiments in sparse reward environments, where tasks are rewarded with 1 at the end of the trajectory if completed and 0 otherwise.
+In this setting, the reward does not accumulate as in our redistribution model.
+As shown in Table 2, RLTF methods do not perform well in sparse reward environments.
+Our proposed method not only outperforms RLTF baselines but also exceeds the performance of sparse reward baselines.
+Table 2: Performance comparison on sparse reward environments over 6 trials with 1e6 time steps for training, presenting average scores and standard deviations.
+Rewards are given with 1 at the end of the trajectory if completed and 0 otherwise.
+The best and comparable methods based on the paired t-test at the significance level $5\%$ are highlighted in boldface.
+Environment SAC IRCR RRD HC Shaping LIRPG RBT(ours)
+ball_in_cup-catch 646.90 773.19 666.18 76.19 847.88 125.50 972.19
+(231.15) (73.06) (33.02) (67.25) (178.01) (21.25) (5.50)
+reacher-hard 531.88 584.76 347.16 7.57 664.49 8.74 735.67
+(127.94) (269.32) (106.56) (3.50) (170.50) (2.36) (144.90)
 5.2 Case Study
@@
 As shown in Fig.
-5, which spans 1000 steps, RBT-predicted rewards, unobservable true rewards, and observable bagged rewards (presented in a uniform format for better visualization) are compared.
+4, which spans 1000 steps, RBT-predicted rewards, unobservable true rewards, and observable bagged rewards (presented in a uniform format for better visualization) are compared.
 The figure indicates that the rewards predicted by the RBT closely match the trends of the true rewards.
 This observation suggest that the RBT is effective at reconstructing true rewards from bagged rewards, despite the coarse nature of the environmental reward signals.
+Figure 5: Ablation study of reward model components across various environments.
+The chart presents mean and standard deviation of rewards over 6 trials with 1e6 timesteps, showcasing the efficacy of the full proposed method relative to its variants without certain features.
+5.3 Ablation Study
+We conducted comparisons to examine the role of RBT’s modules.
+As shown in Fig.
+5, the full RBT model consistently outperforms its variants, indicating a synergistic effect when all components are used together.
+Performance drops significantly when the bidirectional attention mechanism is removed, especially in complex environments like Ant-v2 and HalfCheetah-v2, suggesting its critical role in accurate reward prediction.
+Additionally, we can observe that removing the next state prediction component weakens the reward model’s understanding of environmental dynamics, reducing reward prediction accuracy and hindering policy learning.
+The greatest performance decline occurs when both the next state prediction and bidirectional self-attention mechanism are absent, underscoring their individual and combined importance in building a robust reward model.
 6 Conclusion
-In this paper, we introduce a general learning framework, Learning from Bagged Rewards (RLBR), and make theoretical connections between our learning objectives and traditional MDPs to ensure the justification of our approach.
-Building on this problem, we propose a Transformer-based reward model, the Reward Bag Transformer (RBT), to efficiently redistribute rewards by interpreting contextual information within bags as well as understanding environmental dynamics.
-The efficacy of RBT is demonstrated through extensive experiments, outperforming existing delayed-reward methods in various reward bag scenarios.
-Besides, our case studies highlight the RBT’s ability to effectively reallocate rewards, maintaining fidelity to the original MDP structure.
-While the sum-form bagged reward currently integrated into RLBR addresses most scenarios, as evidenced by previous studies on trajectory feedback (Ren et al., 2021; Arjona-Medina et al., 2019; Zheng et al., 2018), there is value in exploring other forms of reward aggregation.
-Future research will investigate alternative reward structures, such as maximum values or complex combinations of latent rewards, to better capture the nuances of dynamic real-world environments.
-This exploration aims to enhance the adaptability and effectiveness of the RLBR framework in a broader range of applications.
+In this paper, we introduced a novel learning framework called Reinforcement Learning from Bagged Rewards (RLBR).
+To address the challenges posed by this framework, we established theoretical connections between Bagged Reward MDPs (BRMDPs) and MDPs after reward redistribution, providing a solid foundation for our approach.
+Based on this theory, we proposed a reward model, the Reward Bag Transformer (RBT), designed to efficiently redistribute rewards by interpreting contextual information within bags and understanding environmental dynamics.
+The efficacy of RBT was demonstrated through extensive experiments, where it consistently outperformed existing methods across various reward bag scenarios.
+Additionally, our case studies highlighted RBT’s ability to effectively reallocate rewards while maintaining fidelity to the underlying true reward structure.
 Appendix A Broader Impact
-In this work, we introduce the problem of Reinforcement Learning from Bagged Reward (RLBR), and propose the Reward Bag Transformer to address it.
+In this work, we introduce the problem of Reinforcement Learning from Bagged Reward (RLBR), and propose a Transformer-based reward model that uses bidirectional attention to interpret context within a bag and accurately allocate rewards to each instance.
 On the one hand, we recognize that these techniques could raise some potential issues.
@@
 Theorem 1.
-Consider a BRMDP where the bagged reward is defined as the sum of rewards for state-action pairs from the corresponding MDP contained within the same bag.
-In this context, the set of optimal policies for the standard MDP $\Pi$, aligns with that of the BRMDP $\Pi_{\mathrm{B}}$, implying that $\Pi=\Pi_{\mathrm{B}}$.
+Provide a BRMDP where the bagged reward is regarded as the sum of rewards for state-action pairs of a MDP after reward redistribution contained within the same bag.
+In this context, the set of optimal policies for the BRMDP $\Pi_{\mathrm{B}}$ aligns with that of the MDP after reward redistribution $\Pi$, implying that $\Pi=\Pi_{\mathrm{B}}$.
 Proof.
-Let us consider the structure of bagged rewards within BRMDP.
-For a given bag $\mathit{{B}_{i,n_{i}}}$, the bagged reward $R(\mathit{{B}_{i,n_{i}}})$ is the sum of the individual rewards of the state-action pairs contained within it, given by:
-$R(\mathit{{B}_{i,n_{i}}})=\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t}).$
+For a given bag $\mathit{{B}_{i,n_{i}}}$, if the bagged reward $R(\mathit{{B}_{i,n_{i}}})$ is regarded as the sum of the individual rewards of the state-action pairs contained within it, we can get:
+$R(\mathit{{B}_{i,n_{i}}})=\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t}).$
 Over a complete trajectory $\tau$, the cumulative reward in BRMDP can be expressed as the sum of the rewards from all the bags along the trajectory:
-$\sum_{B\in\mathcal{B}_{\tau}}R(B)=\sum_{i\in\mathcal{I}_{\tau}}R(\mathit{{B}_{i,n_{i}}})=\sum_{i\in\mathcal{I}_{\tau}}\bigg{(}\sum_{t=i}^{i+n_{i}-1}r(s_{t},a_{t})\bigg{)}=\sum_{t=0}^{T-1}r(s_{t},a_{t}),$
+$\sum_{B\in\mathcal{B}_{\tau}}R(B)$ $=\sum_{i\in\mathcal{I}_{\tau}}R(\mathit{{B}_{i,n_{i}}})$
+$=\sum_{i\in\mathcal{I}_{\tau}}\bigg{(}\sum_{t=i}^{i+n_{i}-1}\hat{r}(s_{t},a_{t})\bigg{)}=\sum_{t=0}^{T-1}\hat{r}(s_{t},a_{t}),$
 where $\mathcal{I}_{\tau}$ denotes the set of initial timestep indices for $B\in\mathcal{B}_{\tau}$.
-The policy optimization objective in BRMDP, $J_{\mathrm{B}}(\pi)$, aims to maximize the expected sum of bagged rewards along the trajectory from $t=0$, which is equivalent to maximizing the standard cumulative reward in MDP, $J(\pi)$:
+The policy optimization objective in BRMDP, $J_{\mathrm{B}}(\pi)$, aims to maximize the expected sum of bagged rewards along the trajectory from $t=0$, which is equivalent to maximizing the cumulative reward in redistributed reward MDP, $J(\pi)$:
 $J_{\mathrm{B}}(\pi)$ $=\mathbb{E}_{\mathcal{T}(\pi)}\bigg{[}\sum_{B\in\mathcal{B}_{\tau}}R(B)\bigg{|}\mathcal{G}\bigg{]}$
-$=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}r(s_{t},a_{t})\bigg{]}=J(\pi).$
-Given that the expected cumulative rewards for any policy in the BRMDP and MDP frameworks are equivalent, under the condition of infinite exploration or exhaustive sampling within the state-action space, the sets of optimal policies for each framework also coincide, implying that $\Pi=\Pi_{\mathrm{B}}$.
+$=\mathbb{E}_{\pi,P,\mu}\bigg{[}\sum_{t=0}^{T-1}\hat{r}(s_{t},a_{t})\bigg{]}=J(\pi).$
+Given that the expected cumulative rewards for any policy in the original BRMDP and redistributed reward MDP frameworks are equivalent, under the condition of infinite exploration or exhaustive sampling within the state-action space, the sets of optimal policies for each framework also coincide, implying that $\Pi=\Pi_{\mathrm{B}}$.
 ∎
-Appendix C Experiment Settings and Implementation Details
-Benchmarks with Bagged Rewards.
-We introduced a novel problem setting in the suite of MuJoCo and DeepMind Control Suite locomotion benchmark tasks, termed as bagged rewards.
-Our simulations ran on the OpenAI Gym platform Brockman et al.
-(2016) and the DeepMind Control Suite Tassa et al.
-(2018), featuring tasks that stretched over long horizons with a set maximum trajectory length of $T=1000$.
-We used MuJoCo version 2.0 for our simulations, which is available at http://www.mujoco.org/.
-MuJoCo is licensed under a commercial license, and we have adhered to its terms of service and licensing agreements as stated on the official website.
-The DeepMind Control Suite is available under an Apache License 2.0, and we have complied with its terms of use.
-In this setting, the agent received a bagged reward at the end of each collected bag.
-Reward bag experiments of different bag sizes (5, 25, 50, 100, 200, and 500) and trajectory feedback were set up to verify the effectiveness of the method.
-To evaluate the efficacy of proposed method, commonly used trajectory feedback algorithms were adapted to fit the bagged reward setting as baselines.
-In these experiments, each reward bag was treated as an individual trajectory, and these modified algorithms were applied accordingly.
-Additionally, experiments using standard trajectory feedback were conducted to provide a comparative baseline within the unique setting.
-The total episodic feedback was computed at the end of the trajectory and was the sum of the per-step rewards the agent had collected throughout the episode.
-This experiment setting was the same as some previous works for learning from trajectory feedback Gangwani et al.
-(2020); Ren et al.
-(2021).
-Implementation Details and Hyper-parameter Configuration.
 Table 3: Hyper-parameters of RBT.
@@
 Total gradient steps 10000
+Appendix C Experiment Settings and Implementation Details
+Benchmarks with Bagged Rewards.
+We introduced a novel problem setting in the suite of MuJoCo and DeepMind Control Suite locomotion benchmark tasks, termed as bagged rewards.
+Our simulations ran on the OpenAI Gym platform Brockman et al.
+(2016) and the DeepMind Control Suite Tassa et al.
+(2018), featuring tasks that stretched over long horizons with a set maximum trajectory length of $T=1000$.
+We used MuJoCo version 2.0 for our simulations, which is available at http://www.mujoco.org/.
+MuJoCo is licensed under a commercial license, and we have adhered to its terms of service and licensing agreements as stated on the official website.
+The DeepMind Control Suite is available under an Apache License 2.0, and we have complied with its terms of use.
+Reward bag experiments of different bag sizes (5, 25, 50, 100, 200, and 500) and trajectory feedback were set up to verify the effectiveness of the method.
+To evaluate the efficacy of proposed method, commonly used trajectory feedback algorithms were adapted to fit the bagged reward setting as baselines.
+In these experiments, each reward bag was treated as an individual trajectory, and these modified algorithms were applied accordingly.
+Additionally, experiments using standard trajectory feedback were conducted to provide a comparative baseline within the unique setting.
+The total episodic feedback was computed at the end of the trajectory and was the sum of the per-step rewards the agent had collected throughout the episode.
+This experiment setting was the same as some previous works for learning from trajectory feedback Gangwani et al.
+(2020); Ren et al.
+(2021).
+For the sparse reward experiments, the ball_in_cup-catch task provides a sparse reward of 1 when the ball is in the cup and 0 otherwise; in the reacher-hard task, the reward is 1 when the end effector penetrates the target sphere Tassa et al.
+(2018).
+Therefore, we treated the sparse reward as trajectory feedback and conducted experiments where the entire trajectory was considered as a single reward bag in these environments.
+The bagged reward was identical to the sparse reward provided by the environment.
+Implementation Details and Hyper-parameter Configuration.
 In our experiments, the policy optimization module was implemented based on soft actor-critic (SAC) Haarnoja et al.
@@
 (2022) is available in the supplementary material at https://openreview.net/forum?id=nsjkNB2oKsQ.
-To ensure uniformity in the policy optimization process across all methodologies, each was subjected to 1,000,000 training iterations.
-For the proposed method, we initially collated a dataset comprising 10,000 time steps to pre-train the reward model.
-This model then underwent 100 pre-training iterations, a step deemed essential to adequately prepare the reward model before embarking on the principal policy learning phase.
-Following this initial warm-up period, the reward model was trained for 10 iterations after each new trajectory was incorporated.
-Moreover, to systematically gauge performance and progress, evaluations were carried out at intervals of every 5,000 time steps.
-The computational resources for these procedures were NVIDIA GeForce RTX 2080 Ti GPU clusters with 8GB of memory, dedicated to training and evaluating tasks.
 Table 4: Performance comparison across reward bag with various-length configurations over 3 trials.
@@
 (144.24) (682.19) (282.75) (7.54) (134.06) (166.95)
-Appendix D Additional Experimental Results
-This section provides further analysis and insights through additional experiments to complement the main findings presented in the paper.
-Figure 6: Performance comparison adding SAC with reward shaping (SAC-shaping) in the Hopper-v2 and Walker2d-v2 environments, displaying both the mean and standard deviation over 6 trials with different random seeds.
-The experiments are conducted across six specific fixed bag lengths: 5, 25, 50, 100, 200, and 500, as well as with trajectory feedback, labeled as $9$, in each environment.
-Figure 7: Ablation study of reward model components across various environments.
-The chart presents mean and standard deviation of rewards over 6 trials with 1e6 timesteps, showcasing the efficacy of the full proposed method relative to its variants without certain features.
+Figure 6: Comparison of proposed method and sparse reward baselines. The mean and standard deviation are displayed over 6 trials with different random seeds across a total of 1e6 time steps.
+To ensure uniformity in the policy optimization process across all methodologies, each was subjected to 1,000,000 training iterations.
+For the proposed method, we initially collated a dataset comprising 10,000 time steps to pre-train the reward model.
+This model then underwent 100 pre-training iterations, a step deemed essential to adequately prepare the reward model before embarking on the principal policy learning phase.
+Following this initial warm-up period, the reward model was trained for 10 iterations after each new trajectory was incorporated.
+Moreover, to systematically gauge performance and progress, evaluations were carried out at intervals of every 5,000 time steps.
+The computational resources for these procedures were NVIDIA GeForce RTX 2080 Ti GPU clusters with 8GB of memory, dedicated to training and evaluating tasks.
 Figure 7: Rewards comparison and agent states in a trajectory with a bag length of 100 in the Hopper-v2 environment.
 The top graph compares predicted rewards against true rewards and aggregated bagged rewards.
-D.1 Experimental Result of Various-Length Reward Bags
+Figure 8: Learning curves on Ant-v2 with different length of input sequence in training (Seq len) and predict length during relabeling process (Relabel len), based on 3 independent runs with random initialization.
+Within each of the smaller graphs, the curves represent results from experiments with different bag lengths.
+Specifically, there are three bag lengths evaluated: 25, 100, and what is labeled as $9$, which we interpret as a proxy for trajectory feedback.
+The graph highlighted by the red box indicates our chosen parameter setting for the experiment, which is a input sequence length of 100 and a predict length of 500.
+Appendix D Discussion on Sparse Reward
+The sparse reward setting presents significant challenges due to infrequent feedback, making it difficult for agents to effectively explore the environment and discover successful strategies.
+To address this challenge, various methods have been developed to enhance exploration.
+Reward shaping strategies Ng et al.
+(1999); Hu et al.
+(2020); Tambwekar et al.
+(2019) added rewards to actions in a way that guides the agent towards better policies without altering the original reward function.
+Curiosity-driven methods Pathak et al.
+(2017); Sekar et al.
+(2020) encouraged agents to explore the environment by visiting unseen states, potentially solving tasks with sparse rewards.
+Additionally, curriculum learning in RL Florensa et al.
+(2018); Riedmiller et al.
+(2018) involved presenting an agent with a sequence of tasks with gradually increasing complexity, allowing the agent to eventually solve the initially given sparse reward task.
+Due to the formal similarity between sparse rewards and trajectory feedback, both involving a single reward signal after a sequence, these methods can be directly applied to RLTF.
+However, the key difference lies in the nature of the rewards: sparse rewards typically focus on specific key points, making the reward Markovian, whereas RLTF evaluates the entire trajectory, indicating a Non-Markovian reward.
+Furthermore, sparse reward strategies address the exploration-exploitation trade-off inherent in sparse rewards, while RLTF focuses on distributing the evaluation of the entire trajectory across specific state-action pairs to guide learning.
+Thus, while these methods are applicable, they do not address the unique challenges posed by bagged rewards, as they fail to account for the reward structure within the bags.
+Therefore, they are not ideal solutions for RLTF Gangwani et al.
+(2020); Ren et al.
+(2021).
+In Fig.
+6, we present additional results comparing our method with sparse reward baselines.
+Shaping Tessler et al.
+(2019); Hu et al.
+(2020) and LIRPG Zheng et al.
+(2018) are as described in the main paper.
+The Random Network Distillation (RND) Burda et al.
+(2019) method introduces an exploration bonus for deep reinforcement learning by measuring the prediction error of a randomly initialized neural network, which is a widely recognized and utilized approach for addressing sparse reward challenges.
+We include it as a baseline in our comparison of sparse reward methods for a more comprehensive evaluation.
+These results demonstrate that, despite the superficial similarity between sparse rewards and bagged rewards, the internal structure of the rewards differs, leading to suboptimal performance of sparse reward baselines in bagged reward environments.
+Conversely, as shown in the Section Compare with SOTA Methods of the main paper, our method performs well even in sparse reward environments, surpassing baselines specifically designed for sparse reward scenarios.
+This highlights the broad applicability and adaptability of our approach across different reward settings.
+Appendix E Additional Experimental Results
+This section provides further analysis and insights through additional experiments to complement the main findings presented in our main paper.
+E.1 Experimental Result of Various-Length Reward Bags
 In Table 4, we present the experiment results on various-length reward bags.
@@
 However, our RBT method appears to be less sensitive to changes in bag length, maintaining robust performance even when the bag length is equal to the full trajectory.
-This result aligns with the result in Section 5.1.
-D.2 Ablation Study
-We conducted comparisons to examine the role of RBT’s modules.
-As shown in Fig.
-7, the full RBT model consistently outperforms its variants, indicating a synergistic effect when all components are used together.
-Performance drops significantly when the bidirectional attention mechanism is removed, especially in complex environments like Ant-v2 and HalfCheetah-v2, suggesting its critical role in accurate reward prediction.
-Additionally, we can observe that removing the next state prediction component weakens the reward model’s understanding of environmental dynamics, reducing reward prediction accuracy and hindering policy learning.
-The greatest performance decline occurs when both the next state prediction and bidirectional self-attention mechanism are absent, underscoring their individual and combined importance in building a robust reward model.
-D.3 Rewards Comparison in Hopper-v2 with Bagged Rewards
+This result aligns with the result in the Section Compare with SOTA Methods in our main paper.
+E.2 Rewards Comparison in Hopper-v2 with Bagged Rewards
 Fig.
@@
 This analysis underscores that the RBT can adeptly decode the environmental dynamics and the nuanced reward redistribution even under the setting of RLBR.
-D.4 Reward Shaping
-To complement our findings, we conducted additional experiments on reward shaping, following the naive shaping approach described in Hu et al.
-(2020).
-These experiments were performed in the Hopper-v2 and Walker2d-v2 environments due to the availability of task-specific weights from the original paper.
-As shown in the Fig.
-6, the performance of SAC with reward shaping (SAC-shaping) varies with different bag lengths.
-At shorter bag lengths, SAC-shaping tends to perform worse than standard SAC, possibly because the more frequent but less informative rewards add noise to the training process.
-However, as bag length increases, SAC-shaping slightly outperforms SAC, likely because the shaped rewards offer clearer, longer-term signals that assist in learning effective policies when rewards are less frequent.
-Essentially, reward shaping at longer bag lengths creates a more stable and beneficial learning signal for policy improvement.
-The figure demonstrates that while reward shaping improves SAC performance with larger bag sizes, our RBT method consistently outperforms both.
-RBT excels because it not only smooths the rewards but also redistributes them by considering the context within each bag.
-This comprehensive approach allows RBT to more accurately capture environmental complexities, leading to better performance across different bag lengths, indicating its robustness and efficiency in utilizing extended sequences of data.
-D.5 Architecture Sensitivity
+E.3 Architecture Sensitivity
 Fig.
@@
 This analysis underscores the importance of tuning input sequence length during training and prediction length during reward relabeling for optimal performance, while also demonstrating the ability of the RBT model to learn under different parameter settings, showcasing its flexibility and effectiveness in reinforcement learning tasks.
-Figure 9: Learning curves on Ant-v2 with different length of input sequence in training (Seq len) and predict length during relabeling process (Relabel len), based on 3 independent runs with random initialization.
-Within each of the smaller graphs, the curves represent results from experiments with different bag lengths.
-Specifically, there are three bag lengths evaluated: 25, 100, and what is labeled as $9$, which we interpret as a proxy for trajectory feedback.
-The graph highlighted by the red box indicates our chosen parameter setting for the experiment, which is a input sequence length of 100 and a predict length of 500.